### v1

# Microscopic Analysis on LLM players via Social Deduction Game

Abstract
Recent studies have begun developing autonomous game players for social deduction games using large language models (LLMs). When building LLM players, fine-grained evaluations are crucial for addressing weaknesses in game-playing abilities. However, existing studies have often overlooked such assessments. Specifically, we point out two issues with the evaluation methods employed. First, game-playing abilities have typically been assessed through game-level outcomes rather than specific event-level skills; Second, error analyses have lacked structured methodologies. To address these issues, we propose an approach utilizing a variant of the SpyFall game, named SpyGame. We conducted an experiment with four LLMs, analyzing their gameplay behavior in SpyGame both quantitatively and qualitatively. For the quantitative analysis, we introduced eight metrics to resolve the first issue, revealing that these metrics are more effective than existing ones for evaluating the two critical skills: intent identification and camouflage. In the qualitative analysis, we performed thematic analysis to resolve the second issue. This analysis identifies four major categories that affect gameplay of LLMs. Additionally, we demonstrate how these categories complement and support the findings from the quantitative analysis.
Index Terms:
Large language models, Natural language processing, Intelligent agents, Role-playing Games, Autonomous game players, Social deduction games
I Introduction
Conversation plays a pivotal role in social deduction games (SDGs), such as Among Us and Goose Goose Duck. In these games, players are divided into multiple adversarial groups and interact by executing actions through natural language sentences. Unlike other games, these interactions create tricky conversations where explicit statements often contradict the implicit strategies of the corresponding speaker. For example, in Among Us, imposters explicitly and repeatedly claim to be members of the crew. However, this assertion is part of their strategy to conceal their true identity. Consequently, crew members must uncover these hidden intentions to win. As this example illustrates, SDG players must possess the ability to navigate and manage such tricky communications.
Although such abilities are essential, traditional game agents may struggle to handle the tricky conversations occurring in SDGs. Recent studies have adopted large language models (LLMs) because LLMs are capable of generating fluent dialogues and strategic plans. Researchers believe that these capabilities could make LLMs suitable for supporting SDGs. However, most studies have primarily focused on improving gameplay performance rather than investigating whether LLMs can genuinely support these intricate conversations. Specifically, existing studies have two key issues when evaluating SDG-playing agents: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
First, the issue of macroscopic evaluation refers to the tendency of prior studies to evaluate various game-playing abilities using broad game-level outcomes instead of focusing on specific event-level skills. Existing research often relies on simple summative metrics, such as winning rates, which fail to account for the events that occur during gameplay[1, 2]. While these studies acknowledge the importance of certain skills for success in SDGs, they may overlook the need to measure these skills separately with such simple summative metrics. To scrutinize a specific skill, we suggest using microscopic analysis augmented with several specialized metrics. By performing a multifaceted formative analysis leveraging these metrics, one can reveal the true conversational abilities of LLMs in SDG play. In this study, we propose eight microscopic metrics to evaluate two essential skills: (1) Intent identification, the ability to combine explicit clues to uncover hidden information, and (2) Camouflage, which is the ability to disguise oneself from being detected as an opponent.
Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analyses. While certain studies identified abnormal patterns through simple qualitative analysis, the generalizability of their results was limited by their analytical procedures. For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures, while others did not adequately discuss their qualitative analysis methods for deriving generalizable abnormal patterns. To enhance the generalizability of qualitative analysis, we employ a systematic approach known as thematic analysis[3]. Moreover, we aim to draw meaningful insights by correlating the resulting categories with quantitative findings.
We conduct two studies to address these issues. In Study 1 (Section IV), we introduce microscopic metrics and discuss why macroscopic metrics are insufficient for analyzing specific skills. In Study 2 (Section V), we conduct a thematic analysis to derive categories and link these findings to the results of Study 1. In both studies, we let four LLMs play SpyGame, a variant of SpyFall with simplified actions (Section III-A). Consequently, this study makes the following contributions:
  • We demonstrate effectiveness of using microscopic event-level metrics to analyze performance of LLMs in SDGs. Specifically, we show that eight microscopic metrics offer a more detailed assessment than conventional macroscopic metrics in evaluating gameplay skills in SpyGame.
  • We conduct a qualitative analysis to identify categories of abnormal reasoning patterns in LLMs. Through thematic analysis involving multiple LLMs, we discover four main categories and five subcategories, including exposure, role ambiguity, memory distortion, and dissociation.
  • Finally, we validate our quantitative findings by correlating them with the results of the qualitative analysis.
II Related Works
This section reviews the existing literature in two ways. First, we provide a brief overview of LLMs and the historical attempts at developing LLM-based gameplay agents. Then, we point out their two issues on evaluating their agents: macroscopic evaluation and non-systematic analysis.
II-A Existing LLM-based Game Players
Before discussing LLM-based agents in SDGs, we first explain LLMs. An LLM is a language model trained on enormous corpora, excelling in various downstream tasks without requiring task-specific fine-tuning. Researchers in artificial intelligence have reported that LLMs exhibit human-like skills across diverse tasks, including question answering, communication, and planning [4, 5, 6]. For instance, LLMs have been employed to create chatbots that engage in natural, fluent conversations with humans[7]. Therefore, LLMs have been widely adopted in tasks involving natural conversations, such as humor, empathy, and decision-making[8, 9].
Recently, researchers have employed LLMs to design conversational agents for games. Leveraging their natural conversations, studies have reported that LLMs can play simple conversational games at a human level. For example, Brookins and Debacker[10] demonstrated that GPT-3.5 could play the Dictator game[11] and the Prisoner’s Dilemma game[12], displaying preferences for fairness and cooperation that often exceeded those observed in human players.
After confirming the gameplay ability of LLMs in simple conversational games, researchers have turned their attention to a more complex family of games: SDGs. SDGs require a higher level of communication skills like camouflage, which are not essential for playing simpler games. Therefore, researchers have developed autonomous player agents for SDGs, such as Avalon and Werewolf. For instance, Shi et al.[13] designed an LLM-based agent to play Avalon. To build working agents, they incorporated a reflection mechanism[14], which complemented the memory of LLMs with a human-inspired memory mechanism. Similarly, ReCon[15] offers additional hints to assist LLM players in choosing the best action during decision-making. Xu et al.[16] developed an LLM-based Werewolf player using reinforcement learning. They employed a success condition for specific actions as a reward signal to guide the learning procedure. However, these approaches have two key issues when evaluating the performance of their agents, which are discussed in the following section.
II-B Issues in Evaluating LLM Players
A fine-grained evaluation can point out weaknesses of an agent in the gameplay, which is crucial for designing effective LLM players. However, prior studies have typically employed more coarse-grained approaches. Therefore, these studies encounter two main issues in their evaluations: macroscopic quantitative evaluation and non-systematic qualitative analysis.
First, previous studies have leveraged macroscopic metrics for quantitative evaluation. Some researchers have adopted quantitative metrics, whose calculations are solely based on the results of entire game. For example, Xu et al.[2] evaluated LLM-based Werewolf players utilizing the winning rate, total game duration, and total number of camouflage behaviors. Similarly, Light et al.[17] employed winning rates and action success rates to assess gameplay, despite defining several specific gameplay skills, including deductive reasoning, coordination, collaboration, and deception. While these studies successfully measured the overall gameplay ability of an agent, their discussions remained superficial because the evaluation results did not support the existence of specific skills. Specifically, the winning rate (WR) does not provide insights into skills like deception or reasoning skills, as these skills all contribute to victory. Therefore, microscopic metrics are necessary to evaluate specific skills and conduct a multifaceted formative assessment of the gameplay of an LLM agent.
Second, prior studies have employed non-systematic qualitative analyses. Since macroscopic quantitative analysis alone cannot provide sufficient evidence for specific skills, some researchers have adopted qualitative error analysis to assess whether their agents possess certain skills[1, 17, 18]. For instance, Qiao et al.[1] provided a few samples for error analysis without presenting information on frequency of errors. Similarly, Kim and Kim[18] qualitatively analyzed errors made by LLMs in SpyFall. However, these errors are more related to generation error (e.g., contextual or format errors) than to specific gameplay skills. Though these approaches may provide insights into erroneous behaviors in LLM players, the generalizability of the results is questionable since the error may represent a specific but infrequent case. Also, other researchers have adopted language models to conduct quantitative analyses[19, 15]. For instance, Lan et al.[19] used ChatGPT[20] to assess agent gameplay across five aspects, including leadership and persuasion. While this approach can reduce the need for human labor in qualitative analyses, the reliability of these results is questionable due to potential issues with LLMs have reliability issues (e.g., hallucination) in generating answers. Hence, to obtain a comprehensive understanding of the errors made by LLM players, a systematic qualitative analysis of LLM gameplay is necessary.
In summary, the existing literature has two main issues. To address the first issue of macroscopic quantitative measures, we need microscopic measures that can investigate specific skills rather than evaluating overall gameplay ability. To tackle the second issue of non-systematic qualitative analysis, we need to adopt a systematic method for analyzing gameplay logs. In the following sections, especially in Studies 1 and 2, we demonstrate a method for conducting such required analyses using the SDG named SpyGame.
In summary, existing literature has two issues. To address the first issue of macroscopic quantitative measures, we need microscopic measures which can investigate a specific skills instead of whole gameplay ability. Also, to address the second issue of non-systematic qualitative analysis, we need to adopt a systematic way for analyzing gameplay logs. In the following sections, especially in Study 1 and Study 2, we will demonstrate a way to conduct such required analyses, using a SDG called SpyGame.
III Experiment
This section details the process of collecting gameplay logs from LLMs for both quantitative and qualitative analyses. To evaluate performance of vanilla LLM in SDG play, we let LLMs play an SDG through natural language interactions. Section III-A provides an overview of SpyGame, which is based on SpyFall. To collect the gameplay logs of SpyGame, we employed four LLMs, as detailed in Section III-B. The procedure for collecting logs is outlined in Section III-C.
III-A SpyGame
SpyGame is a modified version of SpyFall played by seven players. We altered Spyfall for two key reasons: First, LLMs may have been exposed to the rules and details of Spyfall during their pre-training phase. This exposure can cause LLMs to embed game-related knowledge from pretraining data in their parameters. Therefore, it is likely that these LLMs use this parametric knowledge as a gameplay hint rather than leveraging their inherent abilities. To mitigate this, we renamed the game to SpyGame.
Second, setting a time budget may influence the reproducibility of our analysis. Various factors in the experimental setup (e.g., the hardware used, network status, and model size) can impact inference time. Thus, to enhance reproducibility, we replaced the total time budget of a game with the total number of “turns” within a game.
III-A1 Team formation
Similar to SpyFall, SpyGame features two teams: six citizen players and one spy player. The citizens share a common location, called a hidden location, with each citizen plays a specific role in that location. However, the spy is not given information about the hidden location or the roles of the citizens. In addition, the citizens are unaware of the identity of the spy.
Players must uncover hidden information to win the game. In each turn, one player asks another player a question about the hidden location, and the questioned player must answer. Through these conversations, citizens try to identify the spy. Also, citizens should need to imply their knowledge of the hidden location without revealing too much detail, to avoid being suspected as the spy. Meanwhile, the spy attempts to determine the hidden location by pretending to be a citizen, gathering clues from the utterances of others. The detailed game rules, location-role sets, and prompts used for LLMs are provided in [anonymized for review].
III-A2 Game Flow
Each turn has a designated leader. The turn leader starts the turn by choosing one of two actions: questioning or accusation. If the leader selects the “question” action, they ask another player a question related to the hidden location. The questioned player must then respond based on their assigned role. Alternatively, if the leader opts for the “accusation” action, they pick the player whom the leader most suspects to be the spy and initiate a “Day Vote.” The other players, excluding the accuser and the accused, then vote on whether they agree with the accusation. If all players agree, the game ends immediately. After completing either the “question” or “accusation” action, the questioned or accused player becomes the leader for the next turn.
In every turn after the first, the spy player can publicly reveal their identity and guess the hidden location. Before each turn begins, the system asks the spy player if they are ready to reveal their identity. The game ends immediately when the spy player decides to reveal their identity. Note that in our implementation, instead of directly asking for the decision of the spy, we ask LLMs how certain they are about their guess.
Despite an accusation or an action of the spy, the game may continue until the final turn: Turn 9 in our implementation. At this point, the system initiates the “Final Vote.” All players secretly vote for the player they most suspect of being the spy. Following the vote, the game ends.
At the end of a game, the spy wins if any of the following conditions is met; otherwise, the citizens win:
  • A citizen is eliminated by a Day Vote.
  • The spy publicly reveals their identity and correctly guesses the hidden location.
  • The Final Vote results in multiple spy candidates, with more than one player receiving the most votes.
  • The Final Vote identifies only one spy candidate who is actually a citizen.
III-B Tested Models
We selected four LLMs for this study: GPT-4[4], GPT-3.5-turbo (Chat-GPT), Gemini-pro (1.0)[5], and LlaMA2-70b-chat[6]. We considered three criteria: model size, reasoning performance, and the existence of publicly-available API. Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] to ensure reasoning ability required for SDG play. Previous studies have reported a general performance trend among these four models: GPT-4 leads in performance, followed by Gemini, GPT-3.5, and LlaMA2.
We employed the following input and output procedures to analyze the vanilla performance of the four LLMs: First, as an input prompt, we provided only the minimal information necessary to play SpyGame. Specifically, the LLMs received inputs, including game rules, role descriptions, conversation history, and instructions for generating the next action. Unlike previous studies, we did not provide any additional information (such as hints) that could affect gameplay. Second, for output generation, we used a two-stage process involving free-form generation and JSON formatting. We requested LLMs to decide the next action in free form with their strategy. Then we instructed GPT-4 to rewrite the free-form output into a structured JSON format, which is compatible with our environment. By adopting these procedures, we attempted to minimize unintended effects stemming from additional information.
III-C Data Collection
To assess the gameplay of the spy players, we categorized citizens into two types: strong and weak. Since the game involves two adversarial teams, the performance of the spy can be influenced by the gameplay of the opposing citizens. If the citizen players are strong enough to quickly identify the spy, the spy may not have enough turns to demonstrate their abilities. Thus, we conducted separate experiments with weak citizens versus a spy and strong citizens versus a spy. Here, we used LLaMA2-70B-chat as the weak citizens and GPT-4 as the strong citizens, based on results from other reasoning tasks[21, 22, 4] and our preliminary experiments.
A total of 168 gameplay logs were collected. In this experiment, we adopted seven hidden locations from the original game. For each location, we ran 24 games: 2 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials. Table I lists the statistics of the collected gameplay logs. On average, each game took 5.4 turns, with 23.8 reasoning steps. Here, the reasoning step means a unit of LLM’s reasoning for taking an action.
TABLE I: Average statistics for all players in the collected logs
Turns Reasoning Steps Utterances Words/Utterance
Spy vs. Strong 6.37 27.98 15.65 21.55
Spy vs. Weak 4.52 19.74 10.96 25.24
Total 5.45 23.86 13.31 23.07
IV Multifaceted Quantitative Analysis
Our study aimed to conduct a multifaceted quantitative analysis of LLM performance in SDGs. We designed seven metrics that enable fine-grained analysis of two key gameplay facets: intent identification and camouflage. Compared with previous studies that utilized single-faceted metrics, our approach allows for an analysis of intermediate gameplay behaviors. In this section, we discuss whether existing single-faceted metrics are sufficient for measuring intermediate behaviors. Next, we demonstrate a multifaceted quantitative analysis using the proposed metrics to explore gameplay behavior.
IV-A Insufficiency of Previous Analysis
Previous studies commonly adopted single-faceted analyses to examine SDG gameplay behavior[17, 16, 2, 1]. The most frequently used metrics are the winning rates (WR) and living rounds (LR). These metrics focus on the final state of the game to measure the performance of a player: WR utilizes winner information, while LR uses the number of turns played.
  • Winning Rate (WR) is the percentage of games won by the spy out of the total number of games.
  • Living Round (LR) is the average number of turns played across all games.
IV-A1 Metric result
Tables II (a) and III present the results of the single-faceted metrics when playing against strong and weak citizens, respectively. The WR and LR exhibited different trends in SDG play performance. In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), LLaMA2 (52.38%), and GPT-3.5 (33.33%). A similar trend was observed for weak citizens. However, in terms of LR, LLaMA2 demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71). This trend is similar when examining the performance against weak citizens. Thus, these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies LLaMA2 as the best.
This discrepancy is due to multiple winning conditions specified in the SDG. These conditions allow players to win games in several ways. Therefore, to accurately identify the skills that contribute to winning, it is necessary to consider multiple abilities simultaneously. However, since WR and LR consider only the final state of the game, these metrics cannot examine different abilities separately. For example, in SpyGame, the LR metric indicates how well a spy survives during the game. However, a longer survival duration does not always guarantee victory, leading to potential discrepancies between LR and WR results. Hence, a multifaceted method is essential for a comprehensive analysis of gameplay performance.
TABLE II: Quantitative results for playing against strong citizens
(a) General (b) Intent identification (c) Camouflage
Winning Living Guess Information Notice Caught Vote Vote
Model Rate (%) Round (#) Success (%) Deduction (%) Rate (%) Rate (%) Rate (%) Entropy
Gemini 57.14 6.95 50.00 30.00 33.33 022.22 35.7 0.348
(14) (20) (9) (42)
GPT-3.5 33.33 3.71 27.78 28.57 28.57 007.14 25.0 0.730
(18) (21) (14) (12)
GPT-4 80.95 5.80 81.25 61.90 61.90 025.00 33.3 0.953
(16) (21) (4) (30)
LlaMA2 52.38 9.00 - 00.00 09.52 100.00 45.2 0.634
(0) (19) (10) (126)
Total 55.95 6.365 35.78 30.86 33.33 037.83 40.47 0.642
Metrics whose lower values indicate better results.
TABLE III: Quantitative results for playing against weak citizens
(a) General (b) Intent identification (c) Camouflage
Winning Living Guess Info. Notice Caught Vote Vote
Model Rate (%) Round (#) Success (%) Catching (%) Rate (%) Rate (%) Rate (%) Entropy
Gemini 90.48 5.28 85.71 61.11 090.48 00.00 07.1 0.420
(14) (18) (2) (42)
GPT-3.5 71.43 2.52 71.43 92.86 076.19 $0$
(21) (14) (6) (0)
GPT-4 90.48 2.42 90.48 87.50 100.00 $0$
(21) (16) (2) (0)
LlaMA2 76.19 7.85 60.00 20.00 071.43 60.00 13.5 0.605
(5) (15) (5) (96)
Total 82.14 4.51 80.32 65.07 084.52 20.00 11.59 0.548
Metrics whose lower values indicate better results.
IV-B Intent identification
We designed multiple metrics to enable a multifaceted analysis of intent identification and camouflage. We operationally define Intent identification as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation. Intent identification has two sub-abilities: “information capture” and “information deduction.” Information capture mainly focuses on how well a spy grabs crucial hints that a citizen might inadvertently reveal. Since strong citizens rarely make such mistakes, it is highly challenging to evaluate this sub-ability against them. On the other hand, information deduction (ID) mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information. Since weak citizens usually provide information that directly exposes the hidden location, it is not meaningful to assess this sub-ability against weak citizens. Therefore, metrics should consider whether the spy is effectively utilizing the information provided by these citizens. To evaluate both information capture and deduction, we analyzed outcomes of the guessing by the spy. The following subsections detail the design of the four metrics and their results. The results indicated that GPT-4 was the most powerful spy player, followed by Gemini, GPT-3.5, and LlaMA2.
IV-B1 Metric design
Based on a spy’s guesses during a game, we suggest four metrics for assessing intent identification. Four metrics range from relatively summative to highly formative.
  • Guess Success (GS) measures how much the guessing of a spy contributed to their victory. This metric focuses only on games where the spy attempted to guess, rather than all played games. We calculate GS by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted. A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong intent identification skills.
  • Notice Rate (NR) indicates how well a spy noticed the hidden location across all games. It is calculated by computing the percentage of games in which the spy correctly detected the hidden location over the total number of games played. Unlike GS, which is measured only when the spy decides to reveal their identity publicly, NR can be computed across all games. This is because NR mainly considers whether the spy “notices” the game regardless of whether they decide to guess or not.
  • Information Catching (IC) measures how often the spy successfully leverages exposed information to achieve a victory. Unlike NR, which only assesses the ability of the spy to identify the hidden location, IC evaluates the effectiveness in employing such information to win. We specifically focus on games where weak citizens directly exposed to information about the hidden location directly, instead of considering all played games. We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location guessing from those where exposure occurred. Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing crucial information.
  • Information Deduction (ID) is similar to IC but it applies to games where exposure did not occur. This metric measures how well the spy combined pieces of information to determine the hidden location. To compute ID, we divided the number of games in which the spy successfully guessed the hidden location by the number of games without any exposure. Therefore, a high ID indicates that the spy successfully combined clues from the conversation and utilized them to achieve victory.
In particular, we devised a multifaceted interpretation of intent identification by using both IC and ID metrics simultaneously. By combining the IC and ID with their average values, three combinations were derived. The first combination is the most balanced, where both IC and ID are high. This represents an ideal scenario where the spy successfully identifies the hidden location regardless of the type of citizen. This signifies that the model is good at both sub-abilities: information capture and information deduction. The second combination occurs when only one of the metrics, either IC or ID, is higher than the average. This suggests a lack of proficiency in the corresponding sub-ability. For instance, a model with high IC but low ID demonstrates strong information capture but weak information deduction. The final and least favorable combination for intent identification is when both IC and ID are low. This combination indicates that the spy is unable to infer the hidden location in most games.
Moreover, by combining GS and NR, we can explore whether the spy effectively incorporates the incoming information to maintain their guess. The ideal combination is when both GS and NR are high. This indicates that the spy consistently maintained a correct guess after identifying the hidden location, until the end of the game. In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, he/she manages to maintain their guess once the correct location is identified. Meanwhile, the combination of low GS and high NR implies that although the spy successfully guessed the hidden location, the spy is prone to discarding the guess easily. The worst-case scenario is when both NR and GS are low, indicating a significant failure in both identifying and maintaining the correct hidden location.
IV-B2 Result
Tables II (b) and III (b) present the measurements for intent identification when playing against strong and weak citizens, respectively. The results indicate that GPT-4 has sufficient intent identification ability compared to the other models. GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and LlaMA2 (81.25, 50.0, 27.78, and 0%, respectively). The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and LlaMA2 (9.52%). Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and LlaMA2 (61.90, 30.0, 28.57, and 0%, respectively). In summary, GPT-4 demonstrated the best performance against strong citizens in intent identification. Meanwhile, LlaMA2 showed the lowest performance against strong citizens.
Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and LlaMA2 (90.48, 85.71, 71.43, and 60%, respectively). The NR results displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and LlaMA2 (71.43%). Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and LlaMA2 (20.0%). Similar to the results against strong citizens, GPT-4 showcased good performance against weak citizens.
IV-B3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted intent identification analysis. This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning intent identification.
First, ID and IC are complementary metrics that enable multifaceted analysis of intent identification. As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of intent identification ability. Using these partitions, we can determine the best model possessing sufficient sub-abilities for intent identification; In our analysis, GPT-4 demonstrated high scores in both ID and IC, with its ID being double that of the second-place model. Meanwhile, GPT-3.5 falls within the high-IC region but not the high-ID region. This indicates that while GPT-3.5 struggles to combine hints from the conversation to deduce the hidden location.
Second, by analyzing GS and NR, we can examine how the spy filters incoming information to maintain their guess. As suggested in the metric design, we can use four partitions made by GS and NR. Among the four models, GPT-4 belongs to the first partition and achieves a high score on both metrics. This indicates that GPT-4 can filter useful information, reinforcing the belief of the spy in the hidden location. This could be a hint that GPT-4 has strong information deduction sub-ability. Meanwhile, belonging to different partitions can induce differences in such filtering behaviors. For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, the GS of Gemini was almost double that of GPT-3.5. This difference suggests that Gemini is more selective in using guess actions than GPT-3.5, as evidenced by the fact that Gemini used 14 guess actions regardless of citizens, whereas GPT-3.5 utilized 21 (against weak) and 18 (against strong) guess actions.
Finally, our intent identification metrics enable more detailed analyses of the factors contributing to the scores of simple metrics. Let us consider the simple metric, Win- ning Rate (WR), on Gemini models. Against strong citizens, Gemini showed a similar WR to LlaMA2. Meanwhile, against weak citizens, Gemini exhibited the same WR as GPT-4. However, these observations are somewhat shallow when we consider our intent identification metrics, as the three models display different intent identification abilities. For example, against strong citizens, Gemini had a higher ID score than LLaMA2 in that LLaMA2 completely failed to combine clues about the hidden location. Similarly, against weak citizens, we can conclude that GPT-4 is slightly better than Gemini because of its higher GS and NR scores. Thus, proposed metrics can reveal finer differences between LLMs than existing metrics.
In addition, such a detailed analysis revealed that GPT-3.5 is overly confident in its deduction results. As illustrated in Tables II and III, the spy with GPT-3.5 demonstrated relatively lower NR scores; however, the spy used guessing actions more frequently. That is, the spy rushed into making guesses even when their inference about the hidden location was incorrect. One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as discussed in Section V. We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good intent identification ability.
IV-C Camouflage
In addition to intent identification, we designed four metrics for the multifaceted analysis of camouflage. Here, we operationally define camouflage as the ability to blend in with others during gameplay without revealing one’s true identity. Camouflage encompasses two sub-abilities: “evading suspicion” and “dispersing suspicion.” Evading suspicion mainly considers how well the spy avoids being suspected by others. It is quantified by the number of suspicions directed at the spy. On the other hand, dispersing suspicion primarily considers how difficult the spy makes it for all citizens to converge on a single suspect. It is quantified by the probability distribution of suspicions). To estimate the number and distribution of suspicions, we utilized logs of Day/Final votes. The following subsections describe the design of the four metrics and their results. The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and LlaMA2.
IV-C1 Metric design
Based on day/final votes and accusations during a game, we suggest four metrics for measuring camouflage, ranging from relatively summative to highly formative.
  • Caught Rate (CR) is how frequently the spy is caught by citizens in the Day/Final vote. We compute this by dividing the number of games in which the spy was caught by the number of games lost. Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game. Note that this is a summative metric for the spy’s camouflage performance because it only considers the final state of the game.
  • Accusation Rate (AR) measures how frequently citizens accuse the spy. We measure this by computing the proportion of correct accusations made across all games. Unlike CR, which summarizes all citizens’ thoughts, AR focuses on whether an individual citizen can strongly suspect a true spy since the turn leader is the only player who can make an accusation. Depending on the gameplay, it is possible that there is no accusation during the game; in this case, AR cannot be measured.
  • Vote Rate (VR) is the proportion of suspicions about the spy across all votes that occurred. We measure this by computing the percentage of votes the spy received from citizens across the total number of votes, excluding the spy’s votes. Therefore, a high VR indicates that the spy failed to evade suspicion during the game.
  • Vote Entropy (VE) is how scattered the suspicions about the spy are in the game. We measure this by computing the entropy of the votes that citizens casted for each game and averaging them. Note that in SpyGame, a player’s suspicion cannot affect another player’s suspicion because they cannot have a discussion together. Thus, when a spy successfully blends with citizens, citizens have difficulty narrowing down the number of spy candidates. Thus, the value of VE increases.
We devised a multifaceted interpretation of camouflage by using VR and VE simultaneously. By partitioning VR and VE by their average values, we derived four combinations. First, the most balanced combination was a low VR and high VE. This is an ideal combination because the spy can successfully evade others’ suspicions and lead citizens to discourage each other from narrowing down spy candidates. Second, models in which VR and VE are both low or high lack one of the two sub-abilities. Finally, the worst camouflage combination occurs when VR is high and VE is low. This combination indicates that citizens’ suspicions were concentrated on the spy and that the spy failed to undermine suspicions.
IV-C2 Result
Tables II (c) and III (c) elaborate measurements of camouflage when playing against strong and weak citizens, respectively. We excluded AR from the tables because there was no accusation during gameplay, regardless of the citizen type. Similarly, some values are missing for VR and VE because the spy ended the game early without votes by guessing the hidden location.
The results indicate that the variants of GPT have sufficient camouflage ability compared to other models. Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and LlaMA2 (7.1, 22.2, 25.0, and 100%, respectively). The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and LlaMA2 (45.2%). Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, LlaMA2, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively). Interestingly, GPT-3.5 showed camouflage performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[4].
Against weak citizens, the GPT variants showed good performance compared to others. Because the GPT variants ended the game by guessing, citizens had no chance to find the spy. Thus, the CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models. Other models showed a similar tendency to that against strong citizens. Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than LLaMA2 (60% and 13.5%, respectively). However, regarding VE, Gemini (0.420) performed worse than LLaMA2 (0.605), as shown by the results against strong citizens.
IV-C3 Discussion
In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted camouflage analysis. This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE) and (2) the need for formative camouflage metrics.
First, VR and VE are complementary metrics that enable a multifaceted analysis of camouflage. As suggested above, these two metrics can form three partitions explaining the different levels of camouflage ability. Using these partitions, we can determine the best model for balancing the two sub-abilities of camouflage; the GPT family showed gameplay balanced between VR and VE. These models can evade and disperse suspicion to win the game. In contrast, the worst model in terms of camouflage was LLaMA2. Although LLaMA2 performed better against weak citizens than strong citizens, LLaMA2 player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her. Here, the two sub-abilities are both good or both bad in these two cases. Gemini showed a different behavior: the model was good at evading suspicion but struggled to disperse suspicion. Thus, as illustrated above, VR and VE can provide a wider picture of camouflage behavior.
Second, the proposed metrics provide additional perspectives for investigating camouflage compared to simple summative metrics. In Section IV-A, we observed a discrepancy between two summative metrics, winning rate (WR) and living round (LR). For example, both GPT-3.5 and LLaMA2 showed relatively low WR, but the performance gap regarding LR between them is large. Therefore, using only summative metrics, we can conclude that LLaMA2 outperformed GPT-3.5 in SpyGame. However, our metrics suggested different results; GPT-3.5 is good at camouflage, but LLaMA2 is not. This example shows that the proposed metrics are more suitable for examining camouflage ability than existing summative metrics.
Second, proposed metrics provide additional perspectives on investigating Camouflage, compared to simple summative metrics. In section IV-A, we observed a discrepancy between two summative metrics, winning rate (WR) and living round (LR). For example, both GPT-3.5 and LLaMA2 showed relatively low WR scores, but regarding LR, achieved relatively low and high performance, respectively. So, using summative metrics only, we may conclude that LLaMA2 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggested different results; GPT-3.5 is good at Camouflage, but LLaMA2 is not. This example shows that proposed metrics are more suitable to examine Camouflage ability than existing summative metrics.
Similarly, camouflage metrics are helpful in identifying gameplay problems. For example, both GPT-4 and GPT-3.5 showed relatively low LR scores; however, WR achieved the best and worst scores, respectively. Therefore, using only summative metrics, we can conclude that GPT-4 outperforms GPT-3.5 in SpyGame. However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at camouflage regarding VR and VE.
V Qualitative Analysis
To reconfirm the findings of the quantitative analysis from another perspective, we conducted a thematic analysis. We defined three types of reasoning errors in LLM reasoning while deciding their actions by coding abnormal patterns of 15 randomly selected gameplay logs. As a result, we found that these patterns are related to psychological concepts, including identity confusion, memory distortion, and dissociation[23].
In addition, we found that these concepts are related to the results of the quantitative analysis, especially for caught rate (CR) and guess success (GS). The following subsections illustrate the procedure and results of this qualitative analysis.
V-A Analysis Procedure
Through thematic analysis, we labeled the abnormal patterns of players in the gameplay logs. To enrich our quantitative analysis results, we adopted a bottom-up inductive approach to discover the underlying patterns in the data. The detailed analysis process was as follows. First, we defined the initial labels of the abnormal patterns from 15 randomly selected gameplay logs. Three annotators with computer science backgrounds participated in this process. They read the logs to understand the spy’s gameplay situation when establishing initial labels. To avoid bias towards a specific LLM, we hid the information of the models in the process. Second, we refined these initial labels through an iterative process. Because LLMs perform one or two reasoning steps to decide on an action during the gameplay of SpyGame, a gameplay log is a series of several reasoning steps. So, we used one step as the unit for this analysis. We repeated this process five times between December 29, 2023 and June 1, 2024 until agreement on the identified labels was saturated. The agreement was measured using Fleiss’ Kappa[24], and the final agreement between the annotators ranged from 0.486 to 0.646, which indicates moderate agreement. Finally, using these settled labels, we labeled 506 reasoning steps from 168 games. Consequently, we identified three major categories and five subcategories.
TABLE IV: Frequency of each category found in qualitative analysis
$N$ Role ambiguity Memory distortion Dissociation
Against weak citizens
Gemini 61 1.63 4.91 3.27
GPT-3.5 35 2.85 2.85 8.57
GPT-4 34 0.00 0.00 5.88
LlaMA2 77 9.09 6.49 9.90
Against strong citizens
Gemini 82 3.65 6.09 15.68
GPT-3.5 39 0.00 10.25 23.07
GPT-4 71 4.22 4.22 9.85
LlaMA2 107 9.09 10.28 14.95
The total number of reasoning steps in each condition
V-B Resulting Categories
This section illustrates the results of our thematic analysis. For each category, we explain its definition, report its frequency, and provide some examples. Table IV summarizes the frequencies of all categories.
V-B1 Exposure
Exposure is a case in which a citizen’s utterance literally reveals the hidden location. The Fleiss’ Kappa value for this category was 0.7875. Note that this pattern can only be identified in utterances rather than in reasoning steps. Therefore, we report the frequency of this category based on utterances; among 729 citizen utterances, 125 utterances were labeled as exposure. The following quote provides a representative example:
[The hidden location is an airplane]
LlaMA2 (Citizen, 1st-class passenger):
‘‘What is the maximum speed of an airplane in an emergency situation?’’
We suspect that this exposure pattern is because LLMs may not have sufficient common sense to understand the game rules. In particular, SpyGame has a rule “citizens should not expose the location to the spy.” A human citizen player can notice that a direct mention to other players of the hidden location during a game causes their defeat. However, some LLMs may not have this common sense among players; they simply go through the mentions. Therefore, LLMs believe that they can reveal their information to other citizens because the game rule does not prohibit such behaviors and the citizens already know that fact. Thus, citizen LLMs sometimes publicly reveal their hidden information.
V-B2 Role ambiguity
Role ambiguity[25] is a case in which the spy confuses their own identity with that of a citizen. In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player. Therefore, we divided abnormal patterns of role ambiguity into two sub-categories: team misunderstanding and goal misunderstanding. These two subcategories can occur independently. We present details of the subcategories in the below paragraphs. In summary, Fleiss’ Kappa for this category was 0.563. Among 506 reasoning steps in the data, we identified 21 steps with role ambiguity.
Team misunderstanding
This is a case in which the spy disbelieves that they belong to citizens. Although this can occur in any reasoning step, we provide one example here:
[The spy tries to answer a citizen’s question. Before that, the spy generates its reasoning first.]
Gemini (Spy): Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...
Goal misunderstanding
This is a case in which the spy misbelieves that their goal is the same as that of the citizens. Here is an example:
[The spy tries to decide the next action. After selecting ‘question’ as the next action, the spy generates its reasoning.]
LlaMA2 (Spy): ... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...
We suspect that this category occurs because of the memory limitation of LLMs. To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process. Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input prompt. However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part. For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words. Thus, LLM players can confuse their identities with other characters written in the game rules, triggering role ambiguity.
V-B3 Memory distortion
Memory distortion occurs when a spy uses false or distorted memory to decide their next action. For this category, the Fleiss’ Kappa value was 0.486 among the three annotators. We identified 32 steps with memory distortion among 506 reasoning steps from the data. An example is as follows:
[Previously, Player2 and Player3 did not have any conversation. The spy tries to guess the hidden location.]
GPT-3.5 (Spy): Based on the conversation between Player2 and Player3, it seems that the location is likely ...
We suspect this category is identical to the common “hallucination” phenomenon of LLMs, where a language model generates non-sensical or unfaithful text for the provided source input[26]. In a similar vein, our memory distortion makes the LLM use inappropriate, false, or distorted clues to decide the next action or utterance. These actions, which are affected by hallucination, can hinder the gameplay.
V-B4 Dissociation
Dissociation is a case in which the spy employs their own utterances as evidence of the hidden location because they misunderstand the source of the utterance as other players. The annotators agreed on this category with a Fleiss’ Kappa value of 0.646. In this category, we identified 59 of 506 reasoning steps in the data. An example is as follows:
[Player4, the spy, tries to guess the hidden location.]
LlaMA2 (Spy): Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...
We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols. As previous studies have reported that LLMs suffer from biases induced by proper names[27], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases. Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit. However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes. Thus, LLM players can misinterpret their words as those of others in the conversation log, which triggers dissociation.
V-C Reconfirmation of Quantitative Analysis
Combining the results of the two analyses confirms the findings of the quantitative analysis. Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: camouflage and intent identification.
First, role ambiguity increases the likelihood of a spy’s camouflage failure. When we compared the frequency of role ambiguity and the results of the three camouflage metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the lowest errors in role ambiguity and achieved the highest performance in camouflage metrics. Similarly, LLaMA2 showed more errors while achieving a lower performance than the other models. We suspect that this similarity between the trends is due to the loss of the spy’s need for camouflage. When role ambiguity occurs in the spy’s reasoning steps, the spy confuses their role with that of a citizen. Such confusion leads the spy to make an unusual utterance, which draws suspicion.
Second, memory distortion and dissociation may hinder a spy’s intent identification during the game. When we compared the frequency of memory distortion and dissociation with the results of the four intent identification metrics, the trend of frequency was similar to that of the metrics. For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in intent identification. Similarly, GPT-4 showed fewer errors while achieving a higher performance than the other models. We suspect that this similarity between trends is due to the quality of information that the spy considers during their reasoning process. The spy player should select and combine pieces of information from the conversation to identify the hidden location. When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate because they distorted the information. Similarly, when dissociation occurs, the spy regards the apocryphal information that they previously said as information provided by citizens. Therefore, such misunderstood pieces can be obstacles to reaching the correct answer.
VI Conclusion
To assess whether LLMs can effectively support tricky conversations used in SDGs, we tackled two key issues with previous evaluation methodologies. The first is the reliance on macroscopic quantitative evaluations, which overlook the need for microscopic formative assessment within quantitative analyses. The second is the lack of non-systematic qualitative evaluation, highlighting the necessity for thematic analysis to uncover the logical reasoning behind gameplay. In our demonstration, we chose SpyGame, a variant of SpyFall, as the experimental environment. While human gameplay often incorporates nonverbal and paraverbal cues to identify the hidden intention of others, we adopted the simplest implementation of LLM players to simplify our demonstration. We analyzed the gameplay behaviors using four different LLMs in SpyGame logs. Our analysis was conducted in two ways. First, through quantitative analysis, we employed eight metrics to address the macroscopic evaluation issue. This analysis revealed that our metrics are more effective than existing macroscopic metrics for explaining intent identification and camouflage behavior. Second, in qualitative analysis, we performed thematic analysis to resolve the non-systematic evaluation issue. From the analysis, we discovered four major categories that affect LLMs’ gameplay as a spy. These major categories supported the findings of our quantitative analysis. Despite the limitations of our study, our work underscores the significance of microscopic analysis in evaluating LLMs’ gameplay. In the future, we plan to extend this work to multi-modal LLMs to better capture the differences between human and LLM players. We hope that this study provides valuable insights into the behavior of LLM-based players in SDGs.
VII Acknowledgments
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-01341, Artificial Intelligence Graduate School Program, Chung-Ang University) and the Chung-Ang University Graduate Research Scholarship in 2024.

### v2

@@
-Microscopic Analysis on LLM players via Social Deduction Game
+
 Abstract
-Recent studies have begun developing autonomous game players for social deduction games using large language models (LLMs).
-When building LLM players, fine-grained evaluations are crucial for addressing weaknesses in game-playing abilities.
-However, existing studies have often overlooked such assessments.
-Specifically, we point out two issues with the evaluation methods employed.
-First, game-playing abilities have typically been assessed through game-level outcomes rather than specific event-level skills; Second, error analyses have lacked structured methodologies.
-To address these issues, we propose an approach utilizing a variant of the SpyFall game, named SpyGame.
-We conducted an experiment with four LLMs, analyzing their gameplay behavior in SpyGame both quantitatively and qualitatively.
-For the quantitative analysis, we introduced eight metrics to resolve the first issue, revealing that these metrics are more effective than existing ones for evaluating the two critical skills: intent identification and camouflage.
-In the qualitative analysis, we performed thematic analysis to resolve the second issue.
-This analysis identifies four major categories that affect gameplay of LLMs.
-Additionally, we demonstrate how these categories complement and support the findings from the quantitative analysis.
+Recent studies have investigated whether large language models (LLMs) can support obscure communication that requires specialized skills, such as inferring subtext or doublespeak.
+To conduct the investigation, researchers have used social deduction games (SDGs) as their experimental environment, in which players conceal and infer specific information.
+However, prior work has often overlooked how LLMs should be evaluated in such settings.
+Specifically, we point out two issues with the evaluation methods they employed.
+First, metrics used in prior studies are coarse-grained as they are based on overall game outcomes that often fail to capture event-level behaviors; Second, error analyses have lacked structured methodologies capable of producing insights that meaningfully support evaluation outcomes.
+To address these issues, we propose a macroscopic and systematic approach to the investigation.
+Specifically, we introduce seven fine-grained metrics that resolve the first issue.
+To tackle the second issue, we conducted a thematic analysis and identified four major reasoning failures that undermine LLMs’ performance in obscured communication.
 Index Terms:
 Large language models, Natural language processing, Intelligent agents, Role-playing Games, Autonomous game players, Social deduction games
+[…]
+=-21pt
 I Introduction
-Conversation plays a pivotal role in social deduction games (SDGs), such as Among Us and Goose Goose Duck.
-In these games, players are divided into multiple adversarial groups and interact by executing actions through natural language sentences.
-Unlike other games, these interactions create tricky conversations where explicit statements often contradict the implicit strategies of the corresponding speaker.
-For example, in Among Us, imposters explicitly and repeatedly claim to be members of the crew.
-However, this assertion is part of their strategy to conceal their true identity.
-Consequently, crew members must uncover these hidden intentions to win.
-As this example illustrates, SDG players must possess the ability to navigate and manage such tricky communications.
-Although such abilities are essential, traditional game agents may struggle to handle the tricky conversations occurring in SDGs.
-Recent studies have adopted large language models (LLMs) because LLMs are capable of generating fluent dialogues and strategic plans.
-Researchers believe that these capabilities could make LLMs suitable for supporting SDGs.
-However, most studies have primarily focused on improving gameplay performance rather than investigating whether LLMs can genuinely support these intricate conversations.
-Specifically, existing studies have two key issues when evaluating SDG-playing agents: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
-First, the issue of macroscopic evaluation refers to the tendency of prior studies to evaluate various game-playing abilities using broad game-level outcomes instead of focusing on specific event-level skills.
-Existing research often relies on simple summative metrics, such as winning rates, which fail to account for the events that occur during gameplay[1, 2].
-While these studies acknowledge the importance of certain skills for success in SDGs, they may overlook the need to measure these skills separately with such simple summative metrics.
-To scrutinize a specific skill, we suggest using microscopic analysis augmented with several specialized metrics.
-By performing a multifaceted formative analysis leveraging these metrics, one can reveal the true conversational abilities of LLMs in SDG play.
-In this study, we propose eight microscopic metrics to evaluate two essential skills: (1) Intent identification, the ability to combine explicit clues to uncover hidden information, and (2) Camouflage, which is the ability to disguise oneself from being detected as an opponent.
-Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analyses.
-While certain studies identified abnormal patterns through simple qualitative analysis, the generalizability of their results was limited by their analytical procedures.
-For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures, while others did not adequately discuss their qualitative analysis methods for deriving generalizable abnormal patterns.
-To enhance the generalizability of qualitative analysis, we employ a systematic approach known as thematic analysis[3].
-Moreover, we aim to draw meaningful insights by correlating the resulting categories with quantitative findings.
-We conduct two studies to address these issues.
-In Study 1 (Section IV), we introduce microscopic metrics and discuss why macroscopic metrics are insufficient for analyzing specific skills.
-In Study 2 (Section V), we conduct a thematic analysis to derive categories and link these findings to the results of Study 1.
-In both studies, we let four LLMs play SpyGame, a variant of SpyFall with simplified actions (Section III-A).
+Large language models (LLMs) have shown that they can successfully conduct typical informative communication with humans by leveraging their natural language understanding abilities [1, 2, 3].
+Inspired by these reports, there is a growing interest in adapting LLMs for obscured communication as a next step [4].
+Here, the obscured communication indicates that the intended meanings of messages present in the communication are obscured or made unclear to understand, usually by controlling the spread of information.
+For example, in the real world, an imaginary role-play requires obscure communication to boost immersion in such a role-play.
+Among the required skills, doublespeak is one of the representative skills that enable a participant to achieve an advantageous position in such a play.
+So, since LLMs with doublespeak ability can support the play, researchers have recently focused on investigating such abilities.
+[5]
+Meanwhile, investigating whether LLMs can conduct obscure communication requires a specialized environment.
+As ordinary LLM use cases focus on providing information, obscured communication cannot be tested with such ordinary cases.
+To set a more suitable environment for the examination, researchers have used social deduction games (SDGs) because SDGs force players to use doublespeak to achieve their goals.
+Specifically, they let LLMs play an SDG and collected the game logs, including LLMs’ decisions and conversations.
+After, they devised several metrics using particular game actions for evaluating LLMs’ ability of doublespeak.
+[6] Also, they conducted qualitative analyses to discover erroneous actions or utterances.
+For example, [7] examined whether LLMs can support obscured communication by letting them play a famous SDG, Werewolf.
+However, to thoroughly examine LLMs’ doublespeak abilities, these attempts have two issues: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
+First, the issue of macroscopic evaluation refers to the tendency of prior works to use broad game-level outcomes as their metrics instead of specific event-level actions.
+Existing research often relies on simple metrics, such as win rates, which fail to account for the events that occur during gameplay [6, 7].
+While these studies acknowledge the importance of certain skills for supporting obscured communication, they may overlook the need to measure these skills separately with such granular metrics.
+To scrutinize a specific skill, we suggest using microscopic analysis augmented with several fine-grained metrics.
+Specifically, we propose seven metrics to evaluate two essential skills regarding doublespeak: (1) Subtext inference, the deductive ability to identify doublespeak of others, and (2) Deceptive control, which is the ability to control information to make others confused.
+Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analysis.
+While certain studies identified abnormal patterns through simple qualitative analysis, the reproducibility of their results was limited by their analytical procedures.
+For instance, some studies employed LLMs to analyze data without incorporating human-annotation procedures [8, 9], while others did not adequately discuss their qualitative analysis methods for defining abnormal patterns.
+Also, they simply reported the occurrence tendency of the patterns rather than discussing how they affected their evaluation results [10].
+We believe that conducting qualitative analysis in a systematic way provides both reproducibility and interoperability of experimental results.
+Therefore, we adopt a systematic approach known as thematic analysis [11] to identify recurring reasoning errors that hinder LLMs’ doublespeak.
+By correlating these qualitative patterns with our quantitative metrics, we uncover how specific reasoning failures impact LLMs’ performance in subtext inference and deceptive control.
+We conducted two studies to address these issues.
+In Study 1 (Section IV), we perform a quantitative analysis using our fine-grained metrics.
+The result reveals how coarse-grained metrics have overlooked finer aspects of doublespeak.
+In Study 2 (Section V), we conduct a thematic analysis on collected logs.
+Based on the result, we deduce several categories and link these findings to the results of Study 1.
+In both studies, we let four LLMs play SpyGame, a variant of simple SDG called SpyFall (Section III-A).
 Consequently, this study makes the following contributions:
-  • We demonstrate effectiveness of using microscopic event-level metrics to analyze performance of LLMs in SDGs.
-Specifically, we show that eight microscopic metrics offer a more detailed assessment than conventional macroscopic metrics in evaluating gameplay skills in SpyGame.
-  • We conduct a qualitative analysis to identify categories of abnormal reasoning patterns in LLMs.
-Through thematic analysis involving multiple LLMs, we discover four main categories and five subcategories, including exposure, role ambiguity, memory distortion, and dissociation.
-  • Finally, we validate our quantitative findings by correlating them with the results of the qualitative analysis.
+  • We demonstrate the effectiveness of using microscopic quantitative analysis to assess the doublespeak of LLMs.
+Specifically, using seven fine-grained metrics, we show possibility of studying doublespeak in LLMs’ SDG play.
+  • We identify four categories of abnormal reasoning patterns that harm obscured communication.
+Also, by combining results from the two studies, we discover a correlation between the results of the two analyses.
 II Related Works
 This section reviews the existing literature in two ways.
-First, we provide a brief overview of LLMs and the historical attempts at developing LLM-based gameplay agents.
-Then, we point out their two issues on evaluating their agents: macroscopic evaluation and non-systematic analysis.
-II-A Existing LLM-based Game Players
-Before discussing LLM-based agents in SDGs, we first explain LLMs.
-An LLM is a language model trained on enormous corpora, excelling in various downstream tasks without requiring task-specific fine-tuning.
-Researchers in artificial intelligence have reported that LLMs exhibit human-like skills across diverse tasks, including question answering, communication, and planning [4, 5, 6].
-For instance, LLMs have been employed to create chatbots that engage in natural, fluent conversations with humans[7].
-Therefore, LLMs have been widely adopted in tasks involving natural conversations, such as humor, empathy, and decision-making[8, 9].
-Recently, researchers have employed LLMs to design conversational agents for games.
-Leveraging their natural conversations, studies have reported that LLMs can play simple conversational games at a human level.
-For example, Brookins and Debacker[10] demonstrated that GPT-3.5 could play the Dictator game[11] and the Prisoner’s Dilemma game[12], displaying preferences for fairness and cooperation that often exceeded those observed in human players.
-After confirming the gameplay ability of LLMs in simple conversational games, researchers have turned their attention to a more complex family of games: SDGs.
-SDGs require a higher level of communication skills like camouflage, which are not essential for playing simpler games.
-Therefore, researchers have developed autonomous player agents for SDGs, such as Avalon and Werewolf.
-For instance, Shi et al.[13] designed an LLM-based agent to play Avalon.
-To build working agents, they incorporated a reflection mechanism[14], which complemented the memory of LLMs with a human-inspired memory mechanism.
-Similarly, ReCon[15] offers additional hints to assist LLM players in choosing the best action during decision-making.
-Xu et al.[16] developed an LLM-based Werewolf player using reinforcement learning.
-They employed a success condition for specific actions as a reward signal to guide the learning procedure.
-However, these approaches have two key issues when evaluating the performance of their agents, which are discussed in the following section.
-II-B Issues in Evaluating LLM Players
-A fine-grained evaluation can point out weaknesses of an agent in the gameplay, which is crucial for designing effective LLM players.
-However, prior studies have typically employed more coarse-grained approaches.
-Therefore, these studies encounter two main issues in their evaluations: macroscopic quantitative evaluation and non-systematic qualitative analysis.
-First, previous studies have leveraged macroscopic metrics for quantitative evaluation.
-Some researchers have adopted quantitative metrics, whose calculations are solely based on the results of entire game.
-For example, Xu et al.[2] evaluated LLM-based Werewolf players utilizing the winning rate, total game duration, and total number of camouflage behaviors.
-Similarly, Light et al.[17] employed winning rates and action success rates to assess gameplay, despite defining several specific gameplay skills, including deductive reasoning, coordination, collaboration, and deception.
-While these studies successfully measured the overall gameplay ability of an agent, their discussions remained superficial because the evaluation results did not support the existence of specific skills.
-Specifically, the winning rate (WR) does not provide insights into skills like deception or reasoning skills, as these skills all contribute to victory.
-Therefore, microscopic metrics are necessary to evaluate specific skills and conduct a multifaceted formative assessment of the gameplay of an LLM agent.
-Second, prior studies have employed non-systematic qualitative analyses.
-Since macroscopic quantitative analysis alone cannot provide sufficient evidence for specific skills, some researchers have adopted qualitative error analysis to assess whether their agents possess certain skills[1, 17, 18].
-For instance, Qiao et al.[1] provided a few samples for error analysis without presenting information on frequency of errors.
-Similarly, Kim and Kim[18] qualitatively analyzed errors made by LLMs in SpyFall.
-However, these errors are more related to generation error (e.g., contextual or format errors) than to specific gameplay skills.
-Though these approaches may provide insights into erroneous behaviors in LLM players, the generalizability of the results is questionable since the error may represent a specific but infrequent case.
-Also, other researchers have adopted language models to conduct quantitative analyses[19, 15].
-For instance, Lan et al.[19] used ChatGPT[20] to assess agent gameplay across five aspects, including leadership and persuasion.
-While this approach can reduce the need for human labor in qualitative analyses, the reliability of these results is questionable due to potential issues with LLMs have reliability issues (e.g., hallucination) in generating answers.
-Hence, to obtain a comprehensive understanding of the errors made by LLM players, a systematic qualitative analysis of LLM gameplay is necessary.
-In summary, the existing literature has two main issues.
-To address the first issue of macroscopic quantitative measures, we need microscopic measures that can investigate specific skills rather than evaluating overall gameplay ability.
-To tackle the second issue of non-systematic qualitative analysis, we need to adopt a systematic method for analyzing gameplay logs.
-In the following sections, especially in Studies 1 and 2, we demonstrate a method for conducting such required analyses using the SDG named SpyGame.
-In summary, existing literature has two issues.
-To address the first issue of macroscopic quantitative measures, we need microscopic measures which can investigate a specific skills instead of whole gameplay ability.
-Also, to address the second issue of non-systematic qualitative analysis, we need to adopt a systematic way for analyzing gameplay logs.
-In the following sections, especially in Study 1 and Study 2, we will demonstrate a way to conduct such required analyses, using a SDG called SpyGame.
+First, we outline how recent research has shifted its focus from evaluating LLMs on informative communication tasks to more complex settings that require obscured communication.
+Then, we point out two issues posed in their assessment: macroscopic evaluation and non-systematic analysis.
+II-A Adapting LLMs for Obscured Communication
+LLMs have shown human-level performance in various informative communications.
+For example, researchers widely tested LLMs’ ability to question-answering or summarization [12, 13].
+To accurately and reliably provide answers, LLMs for informative communication attempt to leverage true knowledge and suppress false statements during the generation process [14, 15].
+So, researchers attempt to build an honest and harmless assistant for informative communication.
+Noting that real-world communication includes both informative and obscured communication, researchers have recently turned their interests to obscured communication.
+The two types of communication differ in how each participant views their opponents.
+In informative communication, participants trust their opponents and try to establish a cooperative relationship.
+Meanwhile, participants lack trust in their opponents in obscured communication; thus, they seldom provide valuable information to others and often form adversarial relationships.
+So, participants require different abilities to achieve their goals in such an adversarial situation; skills like doublespeak are needed.
+Specifically, participants should manipulate how they present the information they have or capture subtext from the opponent’s utterances.
+In developing negotiation or debate situations, the ability to discern the opponent’s intention in obscured communication is crucial, as subtext is often used in such communication.
+For example, in the context of applying LLMs to negotiation, prior research has investigated how to measure the ability of an LLM to infer the preferences of the opponent by identifying hidden intentions and meanings in the conversation [16].
+Another example research applied LLMs to debates in order to identify and leverage weaknesses in the opponent’s arguments.
+Here, LLMs’ ability to discern intentions plays a crucial role, helping to accurately recognize hidden intentions or logical weaknesses in the opponent’s claims [17].
+As obscured communication can be naturally seen in games, researchers have recently evaluated relevant skills using social deduction games (SDGs).
+As SDGs force players to utilize obscured communication to win the game, an SDG can be a suitable testbed for assessing or quantifying such communication.
+So, researchers aimed to demonstrate how well their agents can do obscured communication by letting them play SDGs.
+For example, Wang et al.
+[9] proposed an LLM agent capable of playing the Avalon game, introducing mechanisms such as dual contemplation to control conversational information.
+Similarly, Shi et al.
+[18] proposed another LLM agent that can play Avalon.
+Other than these zero-shot attempts, researchers used training methods to improve their agents.
+For instance, Xu et al.
+[19] trained an LLM agent applying reinforcement learning to play the Werewolf game.
+Collectively, these works demonstrate that SDGs offer a structured and adversarial context in which LLMs’ handling of obscured communication can be rigorously assessed.
+II-B Attempts to assess LLMs in obscured communication
+To enhance LLMs’ ability to support obscured communication, it is essential to identify the specific skill areas that require improvement.
+Therefore, most researchers have conducted evaluations in both quantitative and qualitative ways to reveal the current status of LLMs in obscured communications.
+For example, Xu et al.
+[7] and Light et al.
+[20] proposed automated methods for quantitatively evaluating their LLM-based players who play Werewolf or Avalon games.
+Specifically, they adopted particular game conditions, such as win rate or action success rate, that occur during their environmental games as their evaluation metrics.
+However, we suspect that such metrics are too coarse-grained to adequately capture intermediate behaviors during gameplay.
+Specifically, the win rate is only available post hoc, limiting its utility for formative assessment.
+Additionally, the action success rate may fail to account for the multi-player dynamics of SDGs in which performance involves not only task execution but also managing others’ perceptions, such as avoiding suspicion.
+In other words, achieving a higher win rate or action success rate in an SDG does not entail that the model possesses superior competence across all sub-skills involved in obscured communication.
+Other researchers proposed more granular metrics, such as logicality or social behaviors, that can be computed by scoring the reasoning or statements made by LLMs during an Avalon game[9, 8].
+However, such metrics proposed in those studies rely on an auxiliary LLM to qualitatively assess reasonings or statements, rather than being grounded in objective in-game events.
+Therefore, fine-grained quantitative metrics grounded in objective in-game events are essential for evaluating LLMs’ abilities in multi-agent obscured communication, such as subtext inference and deceptive control.
+While quantitative evaluation offers valuable insights into measurable performance, it is often insufficient to fully capture the underlying reasoning errors or abnormal behaviors exhibited by LLMs.
+To address this limitation, researchers have conducted qualitative analyses either as a complement to quantitative evaluation or as a standalone method.
+Kim [10] presented the frequency of misbehavior of LLMs while playing SpyFall to support their assessment of LLMs in the game.
+Quia et al.
+[6] presented some particular cases of LLMs to support their assessment results driven by their evaluation framework using communication games.
+However, the methods used in their qualitative analyses were less structured or not related to quantitative results.
+For example, the categories adopted in [10] related to the generation procedure of LLMs, such as formatting errors, rather than LLMs’ ability to support their obscured environment.
+To enhance the utility of qualitative analysis, it should serve not only to validate the outcomes of quantitative evaluations but also to provide insights into the underlying causes of results.
+In summary, the existing literature has limitations in its quantitative and qualitative analysis methods.
+To discover whether LLMs can support obscured communication, quantitative analysis with fine-grained metrics and validation of qualitative analysis are necessary.
 III Experiment
 This section details the process of collecting gameplay logs from LLMs for both quantitative and qualitative analyses.
-To evaluate performance of vanilla LLM in SDG play, we let LLMs play an SDG through natural language interactions.
+Similar to prior works, we collected gameplay logs of LLMs by letting play SpyGame, which is suitable for observing obscured communication.
 Section III-A provides an overview of SpyGame, which is based on Spyfall.
-To collect the gameplay logs of SpyGame, we employed four LLMs, as detailed in Section III-B.
-The procedure for collecting logs is outlined in Section III-C.
+We employed four LLMs to collect the gameplay logs of SpyGame.
+Section III-B details the tested models and procedure of data collection.
 III-A SpyGame
-SpyGame is a modified version of SpyFall played by seven players.
-We altered Spyfall for two key reasons: First, LLMs may have been exposed to the rules and details of Spyfall during their pre-training phase.
-This exposure can cause LLMs to embed game-related knowledge from pretraining data in their parameters.
-Therefore, it is likely that these LLMs use this parametric knowledge as a gameplay hint rather than leveraging their inherent abilities.
-To mitigate this, we renamed the game to SpyGame.
-Second, setting a time budget may influence the reproducibility of our analysis.
-Various factors in the experimental setup (e.g., the hardware used, network status, and model size) can impact inference time.
-Thus, to enhance reproducibility, we replaced the total time budget of a game with the total number of “turns” within a game.
+Figure 1: SpyGame
+SpyGame is a variation of Spyfall played by seven players [^1] https://github.com/elu-lab/spygame.
+To better assess LLMs’ capacity for obscured communication, we modified the original Spyfall in two ways.
+First, we renamed the game to SpyGame, rather than using Spyfall.
+LLMs may have been exposed to the rules and details of Spyfall during their pre-training phase.
+This exposure can cause LLMs to embed game-related knowledge, such as rules or hints, from pretraining data in their parameters.
+Therefore, it is likely that these LLMs exploit memorized game-specific knowledge from pretraining, rather than demonstrating reasoning-based communicative skills during gameplay.
+So, to avoid such pretrained knowledge, we changed the name of the game to SpyGame.
+Second, we replaced the total time budget of a game with the total number of “turns” within a game.
+Setting a time budget may influence the reproducibility of our analysis.
+Various factors in the experimental setup (e.g., the hardware, network status, and model size) can impact inference time.
+Thus, to enhance reproducibility, we used the number of turns.
 III-A1 Team formation
-Similar to SpyFall, SpyGame features two teams: six citizen players and one spy player.
-The citizens share a common location, called a hidden location, with each citizen plays a specific role in that location.
-However, the spy is not given information about the hidden location or the roles of the citizens.
-In addition, the citizens are unaware of the identity of the spy.
-Players must uncover hidden information to win the game.
-In each turn, one player asks another player a question about the hidden location, and the questioned player must answer.
-Through these conversations, citizens try to identify the spy.
-Also, citizens should need to imply their knowledge of the hidden location without revealing too much detail, to avoid being suspected as the spy.
-Meanwhile, the spy attempts to determine the hidden location by pretending to be a citizen, gathering clues from the utterances of others.
-The detailed game rules, location-role sets, and prompts used for LLMs are provided in [anonymized for review].
+Similar to Spyfall, SpyGame requires two teams: six citizens and one spy.
+The citizens share a common location, called a hidden location, with each citizen playing a specific character in the location.
+For example, let us assume that the location is a school.
+Then, the citizens are assigned to distinct characters, such as a student or a principal.
+Meanwhile, the spy does not have any information about the location or characters.
+Both teams must manage what information they reveal and conceal throughout the game.
+Citizens should protect the information about the location, and the spy should pretend to be a citizen.
+So, to win the game, players must uncover hidden information.
+Simultaneously, all players should avoid being suspected of being a spy.
+Each citizen’s utterance should reveal just enough to signal alignment with other citizens, while minimizing the risk of exposing the hidden location to the spy.
+Meanwhile, the spy should gather subtexts from the conversation and deduce the hidden location.
+This indicates that the abilities required of both the citizen and the spy are centered around concealing and revealing specific information.
+Meanwhile, because SpyGame is adversarial in nature, a substantial imbalance in competence between the evaluated player and their opponents can compromise the validity of the evaluation.
+Moreover, when evaluating a citizen player, the influence of fellow citizens can confound the measurement, making it more difficult to isolate the individual player’s ability.
+For these reasons, we restrict our evaluation to the spy, who operates somewhat independently.
+To assess the performance of spy under varying team conditions, we constructed two scenarios: one with strong citizens and one with weak citizens.
 III-A2 Game Flow
-Each turn has a designated leader.
-The turn leader starts the turn by choosing one of two actions: questioning or accusation.
-If the leader selects the “question” action, they ask another player a question related to the hidden location.
-The questioned player must then respond based on their assigned role.
-Alternatively, if the leader opts for the “accusation” action, they pick the player whom the leader most suspects to be the spy and initiate a “Day Vote.” The other players, excluding the accuser and the accused, then vote on whether they agree with the accusation.
+SpyGame progresses over multiple turns, with each turn consisting of a specific game event.
+Figure 1 illustrates the four key events: (a) questioning and answering, (b) accusation and day vote, (c) guessing, and (d) final vote.
+In each turn, a designated leader initiates the turn by choosing one of two actions: questioning or accusation.
+If the leader chooses to question, they ask another player about the hidden location (Figure 1a).
+Then, the questioned player responds based on their assigned character.
+If the leader opts for an accusation, they point out the player they suspect to be the spy, triggering a Day Vote (Figure 1b).
+All players except the accuser and the accused participate in the vote.
 If all players agree, the game ends immediately.
-After completing either the “question” or “accusation” action, the questioned or accused player becomes the leader for the next turn.
-In every turn after the first, the spy player can publicly reveal their identity and guess the hidden location.
-Before each turn begins, the system asks the spy player if they are ready to reveal their identity.
-The game ends immediately when the spy player decides to reveal their identity.
-Note that in our implementation, instead of directly asking for the decision of the spy, we ask LLMs how certain they are about their guess.
-Despite an accusation or an action of the spy, the game may continue until the final turn: Turn 9 in our implementation.
-At this point, the system initiates the “Final Vote.” All players secretly vote for the player they most suspect of being the spy.
-Following the vote, the game ends.
-At the end of a game, the spy wins if any of the following conditions is met; otherwise, the citizens win:
-  • A citizen is eliminated by a Day Vote.
-  • The spy publicly reveals their identity and correctly guesses the hidden location.
-  • The Final Vote results in multiple spy candidates, with more than one player receiving the most votes.
-  • The Final Vote identifies only one spy candidate who is actually a citizen.
-III-B Tested Models
-We selected four LLMs for this study: GPT-4[4], GPT-3.5-turbo (Chat-GPT), Gemini-pro (1.0)[5], and LlaMA2-70b-chat[6].
-We considered three criteria: model size, reasoning performance, and the existence of publicly-available API.
-Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] to ensure reasoning ability required for SDG play.
-Previous studies have reported a general performance trend among these four models: GPT-4 leads in performance, followed by Gemini, GPT-3.5, and LlaMA2.
-We employed the following input and output procedures to analyze the vanilla performance of the four LLMs: First, as an input prompt, we provided only the minimal information necessary to play SpyGame.
-Specifically, the LLMs received inputs, including game rules, role descriptions, conversation history, and instructions for generating the next action.
-Unlike previous studies, we did not provide any additional information (such as hints) that could affect gameplay.
-Second, for output generation, we used a two-stage process involving free-form generation and JSON formatting.
-We requested LLMs to decide the next action in free form with their strategy.
-Then we instructed GPT-4 to rewrite the free-form output into a structured JSON format, which is compatible with our environment.
-By adopting these procedures, we attempted to minimize unintended effects stemming from additional information.
-III-C Data Collection
-To assess the gameplay of the spy players, we categorized citizens into two types: strong and weak.
-Since the game involves two adversarial teams, the performance of the spy can be influenced by the gameplay of the opposing citizens.
-If the citizen players are strong enough to quickly identify the spy, the spy may not have enough turns to demonstrate their abilities.
-Thus, we conducted separate experiments with weak citizens versus a spy and strong citizens versus a spy.
-Here, we used LLaMA2-70B-chat as the weak citizens and GPT-4 as the strong citizens, based on results from other reasoning tasks[21, 22, 4] and our preliminary experiments.
-A total of 168 gameplay logs were collected.
-In this experiment, we adopted seven hidden locations from the original game.
-For each location, we ran 24 games: 2 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials.
-Table I lists the statistics of the collected gameplay logs.
-On average, each game took 5.4 turns, with 23.8 reasoning steps.
-Here, the reasoning step means a unit of LLM’s reasoning for taking an action.
+After completing either questioning or accusation, the questioned or accused player becomes the leader for the next turn.
+From the second turn, the spy player guesses the hidden location by herself (Figure 1c).
+For each turn, she tells her guess to the system secretly and provides a number indicating how certain she is.
+When the certainty is high enough [^2] In our implementation, we use 9 as the threshold out of 10., the system announces her guess and ends the game.
+Otherwise, the game continues.
+To avoid endless play, we set a final turn [^3] In our implementation, we set the last turn as 9..
+When players finish the final turn, the “Final Vote” begins (Figure 1d).
+All players secretly vote for the most suspicious player in their minds.
+At the end of a game, the spy wins when one of the following conditions is satisfied.
+Otherwise, the citizens win.
+  • All players agreed to eliminate a citizen in a Day Vote.
+  • After the Final Vote, a citizen has the highest vote without a tie.
+  • The spy’s guess is correct when the system announces the spy’s guess.
+III-B Models and Data
+We selected four LLMs for our experiment: GPT-4[1], GPT-3.5-turbo (ChatGPT), Gemini-pro 1.0[3], and Llama2-70b-chat[2].
+As obscured communication requires reasoning on subtexts, we selected large enough models with sufficient reasoning ability on simple tasks.
+Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] in 2023.
+LLMs provide their actions or utterances with their reasoning in natural language.
+So, we employed GPT-4 to sanitize their outputs into a machine-readable format.
+To collect data, we let each tested model play SpyGame as a spy against strong or weak citizens.
+Here, the strong or weak citizen models were fixed based on performance reported in prior studies [21, 22, 1]: GPT-4 for strong and Llama2-70b-chat for weak citizens.
+And, we conducted our experiment repeatedly.
+Specifically, we used seven locations borrowing from the original game.
+For each location, we ran 24 games4442 types of citizens $\times$ 4 spy LLMs $\times$ 3 trials..
+Consequently, we collected a total of 168 gameplay logs.
+Table I lists statistics of collected logs.
+On average, each game took 5.4 turns, with 23.8 reasoning steps which refers to a unit of LLM’s reasoning for taking an action.
 TABLE I: Average statistics for all players in the collected logs
@@
 Total 5.45 23.86 13.31 23.07
-IV Multifaceted Quantitative Analysis
-Our study aimed to conduct a multifaceted quantitative analysis of LLM performance in SDGs.
-We designed seven metrics that enable fine-grained analysis of two key gameplay facets: intent identification and camouflage.
-Compared with previous studies that utilized single-faceted metrics, our approach allows for an analysis of intermediate gameplay behaviors.
-In this section, we discuss whether existing single-faceted metrics are sufficient for measuring intermediate behaviors.
-Next, we demonstrate a multifaceted quantitative analysis using the proposed metrics to explore gameplay behavior.
+IV Microscopic Quantitative Analysis
+Our study aimed to investigate whether LLMs can support obscured communication by conducting a microscopic analysis of their SDG playing.
+We designed seven metrics that enable fine-grained analysis of two key SDG facets: subtext inference and deceptive control.
+Compared with most metrics adopted by previous research, which were designed on game-level outcomes, our approach allows for an analysis of intermediate behaviors during the game.
+Also, our quantitative metrics are fully automated metrics that do not require any manual labeling.
+In this section, we first discuss whether existing simple metrics are sufficient for measuring intermediate behaviors of LLMs.
+Then, we demonstrate our analysis can discover new insights that could not be captured by using current metrics by discussing our quantitative results.
 IV-A Insufficiency of Previous Analysis
-Previous studies commonly adopted single-faceted analyses to examine SDG gameplay behavior[17, 16, 2, 1].
-The most frequently used metrics are the winning rates (WR) and living rounds (LR).
-These metrics focus on the final state of the game to measure the performance of a player: WR utilizes winner information, while LR uses the number of turns played.
-  • Winning Rate (WR) is the percentage of games won by the spy out of the total number of games.
-  • Living Round (LR) is the average number of turns played across all games.
+Previous studies commonly adopted macroscopic analyses to examine SDG gameplay behavior[8, 20, 19, 7, 6].
+The most frequently used metrics are the win rates (WR) and living rounds (LR).
+These metrics evaluate a player’s performance based on final game outcomes: WR is determined by the win/loss result, while LR reflects how long the player remained active during the game.
+  • Win Rate (WR) is the percentage of games won by the spy out of the total number of games.
+  • Living Round (LR) is the average number of turns the spy survived, across all games.
 IV-A1 Metric result
-Tables II (a) and III present the results of the single-faceted metrics when playing against strong and weak citizens, respectively.
+Tables II (a) and III (a) present the results of the coarse-grained metrics when playing against strong and weak citizens, respectively.
 The WR and LR exhibited different trends in SDG play performance.
-In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), LLaMA2 (52.38%), and GPT-3.5 (33.33%).
-A similar trend was observed for weak citizens.
-However, in terms of LR, LLaMA2 demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71).
-This trend is similar when examining the performance against weak citizens.
-Thus, these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies LLaMA2 as the best.
+In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), Llama (52.38%), and GPT-3.5 (33.33%).
+A similar order of performance was observed for against weak citizens.
+However, in terms of LR, Llama demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71).
+This trend is similar when examining the performance playing against weak citizens.
+Note that these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies Llama.
 This discrepancy is due to multiple winning conditions specified in the SDG.
@@
 However, since WR and LR consider only the final state of the game, these metrics cannot examine different abilities separately.
-For example, in SpyGame, the LR metric indicates how well a spy survives during the game.
-However, a longer survival duration does not always guarantee victory, leading to potential discrepancies between LR and WR results.
-Hence, a multifaceted method is essential for a comprehensive analysis of gameplay performance.
+For example, in SpyGame, LR indicates how well a spy survives during the game.
+However, a longer survival does not always guarantee victory, leading to potential discrepancies between LR and WR.
+Hence, a microscopic method is required for a comprehensive analysis of LLM’s performance in various facets.
 TABLE II: Quantitative results for playing against strong citizens
-(a) General (b) Intent identification (c) Camouflage
-Winning Living Guess Information Notice Caught Vote Vote
-Model Rate (%) Round (#) Success (%) Deduction (%) Rate (%) Rate (%) Rate (%) Entropy
-Gemini 57.14 6.95 50.00 30.00 33.33 022.22 35.7 0.348
+(a) General (b) Subtext inference (c) Deceptive Control
+Winning Living Guess Notice Information Information Caught Vote Vote
+Model Rate (%) Round (#) Success (%) Rate (%) Catching Deduction (%) Rate (%) Rate (%) Entropy
+Gemini 57.14 6.95 50.00 $3$ 30.00 022.22 35.7 0.348
 (14) (20) (9) (42)
-GPT-3.5 33.33 3.71 27.78 28.57 28.57 007.14 25.0 0.730
+GPT-3.5 33.33 3.71 27.78 28.57 - 28.57 007.14 25.0 0.730
 (18) (21) (14) (12)
-GPT-4 80.95 5.80 81.25 61.90 61.90 025.00 33.3 0.953
+GPT-4 80.95 5.80 81.25 61.90 - 61.90 025.00 33.3 0.953
 (16) (21) (4) (30)
-LlaMA2 52.38 9.00 - 00.00 09.52 100.00 45.2 0.634
+Llama2 52.38 9.00 - 09.52 - 00.00 100.00 45.2 0.634
 (0) (19) (10) (126)
-Total 55.95 6.365 35.78 30.86 33.33 037.83 40.47 0.642
+Total 55.95 6.365 35.78 $3$ 30.86 037.83 40.47 0.642
 Metrics whose lower values indicate better results.
+Note that Information Catching was not measured playing against strong citizens.
 TABLE III: Quantitative results for playing against weak citizens
-(a) General (b) Intent identification (c) Camouflage
-Winning Living Guess Info. Notice Caught Vote Vote
-Model Rate (%) Round (#) Success (%) Catching (%) Rate (%) Rate (%) Rate (%) Entropy
-Gemini 90.48 5.28 85.71 61.11 090.48 00.00 07.1 0.420
+(a) General (b) Subtext inference (c) Deceptive control
+Winning Living Guess Notice Information Information Caught Vote Vote
+Model Rate (%) Round (#) Success (%) Rate (%) Catching (%) Deduction Rate (%) Rate (%) Entropy
+Gemini 90.48 5.28 85.71 090.48 61.11 - 00.00 07.1 0.420
 (14) (18) (2) (42)
-GPT-3.5 71.43 2.52 71.43 92.86 076.19 $0$
+GPT-3.5 71.43 2.52 71.43 076.19 92.86 - $0$
 (21) (14) (6) (0)
-GPT-4 90.48 2.42 90.48 87.50 100.00 $0$
+GPT-4 90.48 2.42 90.48 100.00 87.50 - $0$
 (21) (16) (2) (0)
-LlaMA2 76.19 7.85 60.00 20.00 071.43 60.00 13.5 0.605
+Llama2 76.19 7.85 60.00 071.43 20.00 - 60.00 13.5 0.605
 (5) (15) (5) (96)
-Total 82.14 4.51 80.32 65.07 084.52 20.00 11.59 0.548
+Total 82.14 4.51 80.32 084.52 65.07 - 20.00 11.59 0.548
 Metrics whose lower values indicate better results.
-IV-B Intent identification
-We designed multiple metrics to enable a multifaceted analysis of intent identification and camouflage.
-We operationally define Intent identification as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation.
-Intent identification has two sub-abilities: “information capture” and “information deduction.” Information capture mainly focuses on how well a spy grabs crucial hints that a citizen might inadvertently reveal.
-Since strong citizens rarely make such mistakes, it is highly challenging to evaluate this sub-ability against them.
-On the other hand, information deduction (ID) mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information.
-Since weak citizens usually provide information that directly exposes the hidden location, it is not meaningful to assess this sub-ability against weak citizens.
-Therefore, metrics should consider whether the spy is effectively utilizing the information provided by these citizens.
-To evaluate both information capture and deduction, we analyzed outcomes of the guessing by the spy.
+Note that Information Deduction was not measured playing against weak citizens.
+IV-B Subtext Inference
+We designed multiple metrics to enable a microscopic analysis of subtext inference.
+We operationally define subtext inference as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation.
+Subtext inference has two sub-abilities: information capture and information deduction.
+Information capture measures whether a spy notices and exploits critical hints inadvertently revealed by citizens.
+Since weak citizens generally leak such information unintentionally, information capture ability becomes essential for the spy to win the game against them.
+Since strong citizens rarely make such mistakes, it is highly challenging to observe information capture of a spy.
+On the other hand, information deduction mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information.
+Since strong citizens generally conceal the key information, information deduction ability becomes essential for the spy to uncover information embodied in a citizen’s utterance.
+To evaluate both information capture and deduction, we analyzed the outcomes of the spy’s guessing.
 The following subsections detail the design of the four metrics and their results.
-The results indicated that GPT-4 was the most powerful spy player, followed by Gemini, GPT-3.5, and LlaMA2.
+The results indicated that GPT-4 was the most powerful spy player, followed by Gemini, GPT-3.5, and Llama.
 IV-B1 Metric design
-Based on a spy’s guesses during a game, we suggest four metrics for assessing intent identification. Four metrics range from relatively summative to highly formative.
+To assess the spy’s ability to identify hidden intent, we propose four metrics based on their guessing behavior. These metrics range from relatively summative to more formative.
   • Guess Success (GS) measures how much the guessing of a spy contributed to their victory.
@@
 We calculate GS by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted.
-A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong intent identification skills.
+A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong subtext inference skills.
   • Notice Rate (NR) indicates how well a spy noticed the hidden location across all games.
@@
 Unlike GS, which is measured only when the spy decides to reveal their identity publicly, NR can be computed across all games.
-This is because NR mainly considers whether the spy “notices” the game regardless of whether they decide to guess or not.
+This is because NR mainly considers whether the spy notices the hidden location, regardless of whether they decide to guess.
   • Information Catching (IC) measures how often the spy successfully leverages exposed information to achieve a victory.
 Unlike NR, which only assesses the ability of the spy to identify the hidden location, IC evaluates the effectiveness in employing such information to win.
-We specifically focus on games where weak citizens directly exposed to information about the hidden location directly, instead of considering all played games.
-We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location guessing from those where exposure occurred.
-Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing crucial information.
+We specifically focus on games where weak citizens directly exposed information about the hidden location instead of considering all the played games.
+Here, the exposure means a case where a citizen’s utterance literally contains the hidden location [^5] Qualitative analysis in Section V-B1 discusses the case..
+We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location from games where exposure occurred.
+Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing key information.
   • Information Deduction (ID) is similar to IC but it applies to games where exposure did not occur.
@@
 Therefore, a high ID indicates that the spy successfully combined clues from the conversation and utilized them to achieve victory.
-In particular, we devised a multifaceted interpretation of intent identification by using both IC and ID metrics simultaneously.
-By combining the IC and ID with their average values, three combinations were derived.
-The first combination is the most balanced, where both IC and ID are high.
-This represents an ideal scenario where the spy successfully identifies the hidden location regardless of the type of citizen.
-This signifies that the model is good at both sub-abilities: information capture and information deduction.
+We provide a multifaceted interpretation of subtext inference by jointly analyzing IC and ID metrics.
+Based on whether each metric is above or below its average, we classify model behavior into three combinations.
+The first combination, when both IC and ID are high, represents an ideal scenario in which the spy successfully identifies the hidden location regardless of the type of citizen.
 The second combination occurs when only one of the metrics, either IC or ID, is higher than the average.
-This suggests a lack of proficiency in the corresponding sub-ability.
+This suggests a relative lack of proficiency in the corresponding sub-ability.
 For instance, a model with high IC but low ID demonstrates strong information capture but weak information deduction.
-The final and least favorable combination for intent identification is when both IC and ID are low.
-This combination indicates that the spy is unable to infer the hidden location in most games.
+The final and least favorable combination for subtext inference is when both IC and ID are low.
+This combination indicates that the spy failed to infer the hidden location in most games.
 Moreover, by combining GS and NR, we can explore whether the spy effectively incorporates the incoming information to maintain their guess.
@@
 This indicates that the spy consistently maintained a correct guess after identifying the hidden location, until the end of the game.
-In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, he/she manages to maintain their guess once the correct location is identified.
+In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, they managed to maintain their guess once the correct location was identified.
 Meanwhile, the combination of low GS and high NR implies that although the spy successfully guessed the hidden location, the spy is prone to discarding the guess easily.
@@
 IV-B2 Result
-Tables II (b) and III (b) present the measurements for intent identification when playing against strong and weak citizens, respectively.
-The results indicate that GPT-4 has sufficient intent identification ability compared to the other models.
-GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and LlaMA2 (81.25, 50.0, 27.78, and 0%, respectively).
-The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and LlaMA2 (9.52%).
-Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and LlaMA2 (61.90, 30.0, 28.57, and 0%, respectively).
-In summary, GPT-4 demonstrated the best performance against strong citizens in intent identification.
-Meanwhile, LlaMA2 showed the lowest performance against strong citizens.
-Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and LlaMA2 (90.48, 85.71, 71.43, and 60%, respectively).
-The NR results displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and LlaMA2 (71.43%).
-Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and LlaMA2 (20.0%).
+Tables II (b) and III (b) present the measurements for subtext inference when playing against strong and weak citizens, respectively.
+The results indicate that GPT-4 has sufficient ability compared to the other models.
+GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (81.25, 50.0, 27.78, and 0%, respectively).
+The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and Llama (9.52%).
+Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and Llama (61.90, 30.0, 28.57, and 0%).
+In summary, GPT-4 demonstrated the best performance against strong citizens in subtext inference.
+Meanwhile, Llama showed the lowest performance against strong citizens.
+Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (90.48, 85.71, 71.43, and 60%, respectively).
+NR displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and Llama (71.43%).
+Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and Llama (20.0%).
 Similar to the results against strong citizens, GPT-4 showcased good performance against weak citizens.
 IV-B3 Discussion
-In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted intent identification analysis.
-This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning intent identification.
-First, ID and IC are complementary metrics that enable multifaceted analysis of intent identification.
-As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of intent identification ability.
-Using these partitions, we can determine the best model possessing sufficient sub-abilities for intent identification; In our analysis, GPT-4 demonstrated high scores in both ID and IC, with its ID being double that of the second-place model.
+In this section, we derive insights using the proposed metrics by conducting a multifaceted analysis.
+This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning subtext inference.
+First, IC and ID are complementary metrics that enable multifaceted analysis of subtext inference.
+As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of subtext inference ability.
+Using these partitions, we can determine the best model possessing sufficient sub-abilities for subtext inference; In our analysis, GPT-4 demonstrated high scores in both IC and ID, with its ID being double that of the second-place model.
 Meanwhile, GPT-3.5 falls within the high-IC region but not the high-ID region.
@@
 Second, by analyzing GS and NR, we can examine how the spy filters incoming information to maintain their guess.
-As suggested in the metric design, we can use four partitions made by GS and NR.
+As suggested in the metric design, we can use four quadrants made by GS and NR.
 Among the four models, GPT-4 belongs to the first partition and achieves a high score on both metrics.
@@
 Meanwhile, belonging to different partitions can induce differences in such filtering behaviors.
-For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, the GS of Gemini was almost double that of GPT-3.5.
+For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, GS of Gemini was almost double that of GPT-3.5.
 This difference suggests that Gemini is more selective in using guess actions than GPT-3.5, as evidenced by the fact that Gemini used 14 guess actions regardless of citizens, whereas GPT-3.5 utilized 21 (against weak) and 18 (against strong) guess actions.
-Finally, our intent identification metrics enable more detailed analyses of the factors contributing to the scores of simple metrics.
-Let us consider the simple metric, Win- ning Rate (WR), on Gemini models.
-Against strong citizens, Gemini showed a similar WR to LlaMA2.
+Finally, our metrics regarding subtext inference enable more detailed analyses of the factors contributing to the scores of simple metrics.
+Let us consider the simple metric, Winning Rate (WR), on Gemini models.
+Against strong citizens, Gemini showed a similar WR to Llama.
 Meanwhile, against weak citizens, Gemini exhibited the same WR as GPT-4.
-However, these observations are somewhat shallow when we consider our intent identification metrics, as the three models display different intent identification abilities.
-For example, against strong citizens, Gemini had a higher ID score than LLaMA2 in that LLaMA2 completely failed to combine clues about the hidden location.
+However, these observations are somewhat shallow when we consider our metrics, as the three models display different subtext inference abilities.
+For example, against strong citizens, Gemini had a higher ID score than Llama in that Llama completely failed to combine clues about the hidden location.
 Similarly, against weak citizens, we can conclude that GPT-4 is slightly better than Gemini because of its higher GS and NR scores.
@@
 That is, the spy rushed into making guesses even when their inference about the hidden location was incorrect.
-One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as discussed in Section V.
-We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good intent identification ability.
-IV-C Camouflage
-In addition to intent identification, we designed four metrics for the multifaceted analysis of camouflage.
-Here, we operationally define camouflage as the ability to blend in with others during gameplay without revealing one’s true identity.
-Camouflage encompasses two sub-abilities: “evading suspicion” and “dispersing suspicion.” Evading suspicion mainly considers how well the spy avoids being suspected by others.
+One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as we will discuss in Section V.
+We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good subtext inference ability.
+IV-C Deceptive Control
+In addition to subtext inference, we designed four metrics for the multifaceted analysis of deceptive control.
+Here, we operationally define deceptive control as the ability to blend in with others by controlling information.
+Deceptive control encompasses two sub-abilities: evading suspicion and dispersing suspicion.
+Evading suspicion mainly considers how well the spy avoids being suspected by others.
 It is quantified by the number of suspicions directed at the spy.
 On the other hand, dispersing suspicion primarily considers how difficult the spy makes it for all citizens to converge on a single suspect.
-It is quantified by the probability distribution of suspicions).
+It is quantified by the probability distribution of suspicions.
 To estimate the number and distribution of suspicions, we utilized logs of Day/Final votes.
 The following subsections describe the design of the four metrics and their results.
-The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and LlaMA2.
+The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and Llama.
 IV-C1 Metric design
-Based on day/final votes and accusations during a game, we suggest four metrics for measuring camouflage, ranging from relatively summative to highly formative.
+Based on day/final votes and accusations during a game, we suggest three metrics for measuring deceptive control, ranging from relatively summative to highly formative.
   • Caught Rate (CR) is how frequently the spy is caught by citizens in the Day/Final vote.
@@
 Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game.
-Note that this is a summative metric for the spy’s camouflage performance because it only considers the final state of the game.
-  • Accusation Rate (AR) measures how frequently citizens accuse the spy.
-We measure this by computing the proportion of correct accusations made across all games.
-Unlike CR, which summarizes all citizens’ thoughts, AR focuses on whether an individual citizen can strongly suspect a true spy since the turn leader is the only player who can make an accusation.
-Depending on the gameplay, it is possible that there is no accusation during the game; in this case, AR cannot be measured.
+Note that this is a summative metric for the spy’s deceptive control ability because it only considers the final state of the game.
   • Vote Rate (VR) is the proportion of suspicions about the spy across all votes that occurred.
@@
 Thus, the value of VE increases.
-We devised a multifaceted interpretation of camouflage by using VR and VE simultaneously.
+We devised a multifaceted interpretation of deceptive control by using VR and VE simultaneously.
 By partitioning VR and VE by their average values, we derived four combinations.
@@
 Second, models in which VR and VE are both low or high lack one of the two sub-abilities.
-Finally, the worst camouflage combination occurs when VR is high and VE is low.
+Finally, the worst combination occurs when VR is high and VE is low.
 This combination indicates that citizens’ suspicions were concentrated on the spy and that the spy failed to undermine suspicions.
 IV-C2 Result
-Tables II (c) and III (c) elaborate measurements of camouflage when playing against strong and weak citizens, respectively.
-We excluded AR from the tables because there was no accusation during gameplay, regardless of the citizen type.
-Similarly, some values are missing for VR and VE because the spy ended the game early without votes by guessing the hidden location.
-The results indicate that the variants of GPT have sufficient camouflage ability compared to other models.
-Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and LlaMA2 (7.1, 22.2, 25.0, and 100%, respectively).
-The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and LlaMA2 (45.2%).
-Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, LlaMA2, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively).
-Interestingly, GPT-3.5 showed camouflage performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[4].
-Against weak citizens, the GPT variants showed good performance compared to others.
-Because the GPT variants ended the game by guessing, citizens had no chance to find the spy.
-Thus, the CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models.
-Other models showed a similar tendency to that against strong citizens.
-Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than LLaMA2 (60% and 13.5%, respectively).
-However, regarding VE, Gemini (0.420) performed worse than LLaMA2 (0.605), as shown by the results against strong citizens.
+Tables II (c) and III (c) elaborate measurements of deceptive control when playing against strong and weak citizens, respectively.
+Note that some missing values for VR and VE indicate that the spy guessed the hidden location for all games, and no votes were conducted.
+The results indicate that the variants of GPT have sufficient ability compared to other models.
+Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and Llama (7.1, 22.2, 25.0, and 100%, respectively).
+The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and Llama (45.2%).
+Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, Llama, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively).
+Interestingly, GPT-3.5 showed deceptive control performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[1].
+Against weak citizens, GPT variants showed good performance compared to others.
+Because GPT variants ended the game by guessing, citizens had no chance to find the spy.
+Thus, CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models.
+Other models showed a tendency similar to that against strong citizens.
+Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than Llama (60% and 13.5%).
+However, regarding VE, Gemini (0.420) performed worse than Llama (0.605), as shown by results against strong citizens.
 IV-C3 Discussion
-In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted camouflage analysis.
-This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE) and (2) the need for formative camouflage metrics.
-First, VR and VE are complementary metrics that enable a multifaceted analysis of camouflage.
-As suggested above, these two metrics can form three partitions explaining the different levels of camouflage ability.
-Using these partitions, we can determine the best model for balancing the two sub-abilities of camouflage; the GPT family showed gameplay balanced between VR and VE.
+In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted analysis of deceptive control.
+This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE), and (2) additional perspectives captured by adopting formative metrics.
+First, VR and VE are complementary metrics that enable a multifaceted analysis of deceptive control.
+As suggested above, these two metrics can form three partitions explaining the different levels of capacity regarding deceptive control.
+Using these partitions, we can determine the best model for balancing two sub-abilities of deceptive control; GPT family showed gameplay balanced between VR and VE.
 These models can evade and disperse suspicion to win the game.
-In contrast, the worst model in terms of camouflage was LLaMA2.
-Although LLaMA2 performed better against weak citizens than strong citizens, LLaMA2 player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her.
-Here, the two sub-abilities are both good or both bad in these two cases.
+In contrast, the worst model in terms of deceptive control was Llama.
+Although Llama performed better against weak citizens than strong citizens, Llama player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her.
+So, the two sub-abilities are both good or both bad in these two cases.
 Gemini showed a different behavior: the model was good at evading suspicion but struggled to disperse suspicion.
-Thus, as illustrated above, VR and VE can provide a wider picture of camouflage behavior.
-Second, the proposed metrics provide additional perspectives for investigating camouflage compared to simple summative metrics.
-In Section IV-A, we observed a discrepancy between two summative metrics, winning rate (WR) and living round (LR).
-For example, both GPT-3.5 and LLaMA2 showed relatively low WR, but the performance gap regarding LR between them is large.
-Therefore, using only summative metrics, we can conclude that LLaMA2 outperformed GPT-3.5 in SpyGame.
-However, our metrics suggested different results; GPT-3.5 is good at camouflage, but LLaMA2 is not.
-This example shows that the proposed metrics are more suitable for examining camouflage ability than existing summative metrics.
-Second, proposed metrics provide additional perspectives on investigating Camouflage, compared to simple summative metrics.
-In section IV-A, we observed a discrepancy between two summative metrics, winning rate (WR) and living round (LR).
-For example, both GPT-3.5 and LLaMA2 showed relatively low WR scores, but regarding LR, achieved relatively low and high performance, respectively.
-So, using summative metrics only, we may conclude that LLaMA2 outperforms GPT-3.5 in SpyGame.
-However, the proposed metrics suggested different results; GPT-3.5 is good at Camouflage, but LLaMA2 is not.
-This example shows that proposed metrics are more suitable to examine Camouflage ability than existing summative metrics.
-Similarly, camouflage metrics are helpful in identifying gameplay problems.
+Thus, as illustrated above, VR and VE can provide a wider picture of deceptive control behavior.
+Second, proposed metrics provide additional perspectives on investigating deceptive control, compared to simple summative metrics.
+In section IV-A, we observed a discrepancy between two summative metrics, win rate (WR) and living round (LR).
+For example, both GPT-3.5 and Llama showed relatively low WR scores, but regarding LR, they achieved relatively low and high performance, respectively.
+So, using summative metrics only, we may conclude that Llama outperforms GPT-3.5 in SpyGame.
+However, the proposed metrics suggested different results; GPT-3.5 is good at controlling information, but Llama is not.
+This example shows that our proposed metrics are more suitable for examining deceptive control ability than existing summative metrics.
+Similarly, our proposed metrics are helpful in identifying gameplay problems.
 For example, both GPT-4 and GPT-3.5 showed relatively low LR scores; however, WR achieved the best and worst scores, respectively.
 Therefore, using only summative metrics, we can conclude that GPT-4 outperforms GPT-3.5 in SpyGame.
-However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at camouflage regarding VR and VE.
+However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at deceptive control regarding VR and VE.
 V Qualitative Analysis
-To reconfirm the findings of the quantitative analysis from another perspective, we conducted a thematic analysis.
+To reconfirm the findings of quantitative analysis from another perspective, we conducted a thematic analysis.
 We defined three types of reasoning errors in LLM reasoning while deciding their actions by coding abnormal patterns of 15 randomly selected gameplay logs.
@@
 TABLE IV: Frequency of each category found in qualitative analysis
-$N$ Role ambiguity Memory distortion Dissociation
+$N$* Character ambiguity Memory distortion Dissociation
 Against weak citizens
@@
 Llama2 107 9.09 10.28 14.95
-The total number of reasoning steps in each condition
+* The total number of reasoning steps in each condition
 V-B Resulting Categories
@@
 Note that this pattern can only be identified in utterances rather than in reasoning steps.
-Therefore, we report the frequency of this category based on utterances; among 729 citizen utterances, 125 utterances were labeled as exposure.
+Therefore, we report the frequency of this category based on utterances; among 729 citizen utterances, 125 were labeled as exposure.
 The following quote provides a representative example:
@@
 ‘‘What is the maximum speed of an airplane in an emergency situation?’’
-We suspect that this exposure pattern is because LLMs may not have sufficient common sense to understand the game rules.
+We suspect that LLMs may not have sufficient common sense to understand the game rules, causing this pattern.
 In particular, SpyGame has a rule “citizens should not expose the location to the spy.” A human citizen player can notice that a direct mention to other players of the hidden location during a game causes their defeat.
-However, some LLMs may not have this common sense among players; they simply go through the mentions.
-Therefore, LLMs believe that they can reveal their information to other citizens because the game rule does not prohibit such behaviors and the citizens already know that fact.
+However, some LLMs may not have this common sense among human players; they simply go through the mentions.
+Therefore, LLMs believe that they can reveal their information to other citizens because the game rule does not algorithmically prohibit such behaviors and the citizens already know that fact.
 Thus, citizen LLMs sometimes publicly reveal their hidden information.
-V-B2 Role ambiguity
-Role ambiguity[25] is a case in which the spy confuses their own identity with that of a citizen.
-In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player.
-Therefore, we divided abnormal patterns of role ambiguity into two sub-categories: team misunderstanding and goal misunderstanding.
-These two subcategories can occur independently.
-We present details of the subcategories in the below paragraphs.
-In summary, Fleiss’ Kappa for this category was 0.563.
-Among 506 reasoning steps in the data, we identified 21 steps with role ambiguity.
-Team misunderstanding
-This is a case in which the spy disbelieves that they belong to citizens. Although this can occur in any reasoning step, we provide one example here:
-[The spy tries to answer a citizen’s question. Before that, the spy generates its reasoning first.]
-Gemini (Spy): Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...
-Goal misunderstanding
-This is a case in which the spy misbelieves that their goal is the same as that of the citizens. Here is an example:
-[The spy tries to decide the next action. After selecting ‘question’ as the next action, the spy generates its reasoning.]
-LlaMA2 (Spy): ... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...
-We suspect that this category occurs because of the memory limitation of LLMs.
-To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process.
-Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input prompt.
-However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part.
-For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words.
-Thus, LLM players can confuse their identities with other characters written in the game rules, triggering role ambiguity.
+V-B2 Dissociation
+Dissociation is a case in which the spy employs their own utterances as evidence of the hidden location because they misunderstand the source of the utterance as other players.
+The annotators agreed on this category with a Fleiss’ Kappa value of 0.646.
+In this category, we identified 59 of 506 reasoning steps in the data.
+An example is as follows:
+[Player4, the spy, tries to guess the hidden location.]
+Llama2 (Spy): Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...
+We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols.
+As previous studies have reported that LLMs suffer from biases induced by proper names[25], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases.
+Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit.
+However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes [26].
+Thus, LLM players can misinterpret their words as those of others in the log, which triggers dissociation.
 V-B3 Memory distortion
@@
 These actions, which are affected by hallucination, can hinder the gameplay.
-V-B4 Dissociation
-Dissociation is a case in which the spy employs their own utterances as evidence of the hidden location because they misunderstand the source of the utterance as other players.
-The annotators agreed on this category with a Fleiss’ Kappa value of 0.646.
-In this category, we identified 59 of 506 reasoning steps in the data.
-An example is as follows:
-[Player4, the spy, tries to guess the hidden location.]
-LlaMA2 (Spy): Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...
-We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols.
-As previous studies have reported that LLMs suffer from biases induced by proper names[27], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases.
-Therefore, when an LLM attempts to collect clues from the conversation log, it has to distinguish the source of each utterance using a single difference between player names: a single digit.
-However, previous studies have reported that LLMs struggle to recognize different digit tokens during their computation processes.
-Thus, LLM players can misinterpret their words as those of others in the conversation log, which triggers dissociation.
+V-B4 Character ambiguity
+Character ambiguity is a case in which the spy confuses their own identity with that of a citizen.
+This category is named after a similar psychological concept [28].
+In an SDG, understanding a player’s identity is related to two major concepts affecting gameplay: (1) the team of the player and (2) the goal of the player.
+Therefore, we divided abnormal patterns of character ambiguity into two sub-categories: team misunderstanding and goal misunderstanding.
+These two subcategories can occur independently.
+We present details of the subcategories in the below paragraphs.
+In summary, Fleiss’ Kappa for this category was 0.563.
+Among 506 reasoning steps in the data, we identified 21 steps with character ambiguity.
+Team misunderstanding
+This is a case in which the spy disbelieves that they belong to citizens. Although this can occur in any reasoning step, we provide one example here:
+[The spy tries to answer a citizen’s question. Before that, the spy generates its reasoning first.]
+Gemini (Spy): Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...
+Goal misunderstanding
+This is a case in which the spy misbelieves that their goal is the same as that of the citizens. Here is an example:
+[The spy tries to decide the next action. After selecting ‘question’ as the next action, the spy generates its reasoning.]
+Llama2 (Spy): ... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...
+We suspect that this category occurs is due to the memory limitation of LLMs.
+To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process.
+Since we provided the game rule, identity of a player, and conversation logs together as an input prompt, the LLM needs to distinguish information about the identity from the input.
+However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part.
+For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words.
+Thus, LLM players can confuse their identities with other characters written in the game rules, triggering character ambiguity.
 V-C Reconfirmation of Quantitative Analysis
 Combining the results of the two analyses confirms the findings of the quantitative analysis.
-Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: camouflage and intent identification.
-First, role ambiguity increases the likelihood of a spy’s camouflage failure.
-When we compared the frequency of role ambiguity and the results of the three camouflage metrics, the trend of frequency was similar to that of the metrics.
-For example, against strong citizens, GPT-3.5 showed the lowest errors in role ambiguity and achieved the highest performance in camouflage metrics.
-Similarly, LLaMA2 showed more errors while achieving a lower performance than the other models.
-We suspect that this similarity between the trends is due to the loss of the spy’s need for camouflage.
-When role ambiguity occurs in the spy’s reasoning steps, the spy confuses their role with that of a citizen.
-Such confusion leads the spy to make an unusual utterance, which draws suspicion.
-Second, memory distortion and dissociation may hinder a spy’s intent identification during the game.
-When we compared the frequency of memory distortion and dissociation with the results of the four intent identification metrics, the trend of frequency was similar to that of the metrics.
-For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in intent identification.
+Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: subtext inference and deceptive control.
+First, memory distortion and dissociation may hinder a spy from inferring subtext during the game.
+When we compared the frequency of memory distortion and dissociation with the results of the four subtext inference metrics, the trend of frequency was similar to that of the metrics.
+For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in subtext inference.
 Similarly, GPT-4 showed fewer errors while achieving a higher performance than the other models.
@@
 The spy player should select and combine pieces of information from the conversation to identify the hidden location.
-When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate because they distorted the information.
-Similarly, when dissociation occurs, the spy regards the apocryphal information that they previously said as information provided by citizens.
+When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate.
+Similarly, when dissociation occurs, the spy regards apocryphal information that they previously said as information provided by citizens.
 Therefore, such misunderstood pieces can be obstacles to reaching the correct answer.
+Second, character ambiguity increases the likelihood of a spy’s failure of deceptive control.
+When we compared the frequency of character ambiguity and the results of the three metrics regarding deceptive control, the trend of frequency was similar to that of the metrics.
+For example, against strong citizens, GPT-3.5 showed the lowest errors in character ambiguity and achieved the highest performance in such metrics.
+Also, Llama showed more errors while achieving a lower performance than the other models.
+We suspect that this similarity between the trends is due to the loss of the spy’s need for deception.
+When character ambiguity occurs in the spy’s reasoning steps, the spy confuses their character with that of a citizen.
+Such confusion leads the spy to make an unusual utterance, which draws suspicion.
 VI Conclusion
-To assess whether LLMs can effectively support tricky conversations used in SDGs, we tackled two key issues with previous evaluation methodologies.
-The first is the reliance on macroscopic quantitative evaluations, which overlook the need for microscopic formative assessment within quantitative analyses.
-The second is the lack of non-systematic qualitative evaluation, highlighting the necessity for thematic analysis to uncover the logical reasoning behind gameplay.
-In our demonstration, we chose SpyGame, a variant of SpyFall, as the experimental environment.
-While human gameplay often incorporates nonverbal and paraverbal cues to identify the hidden intention of others, we adopted the simplest implementation of LLM players to simplify our demonstration.
-We analyzed the gameplay behaviors using four different LLMs in SpyGame logs.
-Our analysis was conducted in two ways.
-First, through quantitative analysis, we employed eight metrics to address the macroscopic evaluation issue.
-This analysis revealed that our metrics are more effective than existing macroscopic metrics for explaining intent identification and camouflage behavior.
-Second, in qualitative analysis, we performed thematic analysis to resolve the non-systematic evaluation issue.
-From the analysis, we discovered four major categories that affect LLMs’ gameplay as a spy.
-These major categories supported the findings of our quantitative analysis.
-Despite the limitations of our study, our work underscores the significance of microscopic analysis in evaluating LLMs’ gameplay.
-In the future, we plan to extend this work to multi-modal LLMs to better capture the differences between human and LLM players.
-We hope that this study provides valuable insights into the behavior of LLM-based players in SDGs.
-VII Acknowledgments
-This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2021-0-01341, Artificial Intelligence Graduate School Program, Chung-Ang University) and the Chung-Ang University Graduate Research Scholarship in 2024.
+To assess whether LLMs can effectively support obscured communication, we addressed two key limitations in prior evaluation methodologies.
+The first is the reliance on coarse-grained quantitative metrics, which fail to capture fine-grained communicative behaviors.
+The second is the lack of structured qualitative analyses that can contextualize evaluation results or uncover the reasoning processes underlying model behavior.
+To investigate these limitations in a controlled setting, we conducted experiments using SpyGame, a variant of Spyfall, and collected gameplay logs from four LLMs.
+We analyzed these logs through both quantitative and qualitative methods.
+First, we proposed eight fine-grained metrics designed to evaluate two core abilities: subtext inference and deceptive control.
+These metrics demonstrated more effective in identifying subtle behavioral differences between models.
+Second, we performed a thematic analysis to identify reasoning failures not captured by quantitative metrics.
+This analysis revealed four major categories of failure that align with the patterns observed in the quantitative evaluation, offering complementary insights.
+Our findings highlight the importance of microscopic, multi-perspective evaluation for understanding how LLMs perform in obscured communicative settings.
+We hope this study contributes to a deeper understanding on LLMs in complex social interactions.

### v3

@@
 Abstract
-Recent studies have investigated whether large language models (LLMs) can support obscure communication that requires specialized skills, such as inferring subtext or doublespeak.
+Recent studies have investigated whether large language models (LLMs) can support obscured communication, which is characterized by core aspects such as inferring subtext and evading suspicions.
 To conduct the investigation, researchers have used social deduction games (SDGs) as their experimental environment, in which players conceal and infer specific information.
 However, prior work has often overlooked how LLMs should be evaluated in such settings.
-Specifically, we point out two issues with the evaluation methods they employed.
+Specifically, we point out two limitations with the evaluation methods they employed.
 First, metrics used in prior studies are coarse-grained as they are based on overall game outcomes that often fail to capture event-level behaviors; Second, error analyses have lacked structured methodologies capable of producing insights that meaningfully support evaluation outcomes.
-To address these issues, we propose a macroscopic and systematic approach to the investigation.
-Specifically, we introduce seven fine-grained metrics that resolve the first issue.
+To address these limitations, we propose a microscopic and systematic approach to the investigation.
+Specifically, we introduce six fine-grained metrics that resolve the first issue.
 To tackle the second issue, we conducted a thematic analysis and identified four major reasoning failures that undermine LLMs’ performance in obscured communication.
@@
 Among the required skills, doublespeak is one of the representative skills that enable a participant to achieve an advantageous position in such a play.
-So, since LLMs with doublespeak ability can support the play, researchers have recently focused on investigating such abilities.
-[5]
-Meanwhile, investigating whether LLMs can conduct obscure communication requires a specialized environment.
-As ordinary LLM use cases focus on providing information, obscured communication cannot be tested with such ordinary cases.
+So, since LLMs with doublespeak ability can support the play, researchers have recently focused on investigating such abilities [5].
+In this context, evaluating LLMs’ ability to support obscured communication is challenging because it requires assessing not just informativeness or task completion, but also the ability to strategically conceal sensitive information and infer hidden meanings from indirect cues.
+Unlike cooperative dialogue, where participants work toward shared goals, obscured communication is inherently adversarial, making it a significantly more complex and nuanced form of interaction.
 To set a more suitable environment for the examination, researchers have used social deduction games (SDGs) because SDGs force players to use doublespeak to achieve their goals.
 Specifically, they let LLMs play an SDG and collected the game logs, including LLMs’ decisions and conversations.
-After, they devised several metrics using particular game actions for evaluating LLMs’ ability of doublespeak.
-[6] Also, they conducted qualitative analyses to discover erroneous actions or utterances.
+After, they devised several metrics using particular game actions for evaluating LLMs’ ability of doublespeak [6].
+Also, they conducted qualitative analyses to discover erroneous actions or utterances.
 For example, [7] examined whether LLMs can support obscured communication by letting them play a famous SDG, Werewolf.
-However, to thoroughly examine LLMs’ doublespeak abilities, these attempts have two issues: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
-First, the issue of macroscopic evaluation refers to the tendency of prior works to use broad game-level outcomes as their metrics instead of specific event-level actions.
-Existing research often relies on simple metrics, such as win rates, which fail to account for the events that occur during gameplay [6, 7].
-While these studies acknowledge the importance of certain skills for supporting obscured communication, they may overlook the need to measure these skills separately with such granular metrics.
-To scrutinize a specific skill, we suggest using microscopic analysis augmented with several fine-grained metrics.
-Specifically, we propose seven metrics to evaluate two essential skills regarding doublespeak: (1) Subtext inference, the deductive ability to identify doublespeak of others, and (2) Deceptive control, which is the ability to control information to make others confused.
+However, we question whether these prior attempts are sufficient to thoroughly assess whether LLMs can effectively support obscured communication due to their two limitations: (1) Macroscopic quantitative evaluation and (2) Non-systematic qualitative analysis.
+First, the issue of macroscopic evaluation refers to the tendency of prior works to use broad game-level outcomes as their metrics, rather than specific event-level actions.
+Existing research often relies on simple metrics, such as win rate, which fail to account for the events that occur during gameplay [6, 7].
+While these studies acknowledge the importance of specific aspects for supporting obscured communication, they may overlook the need to measure these aspects separately with such granular metrics.
+This raises our first research question (RQ1): Does a high score based on coarse-grained metrics imply that an LLM supports all the key aspects?
+To address this question, we propose six metrics to evaluate two core aspects that characterize obscured communication: (1) Subtext inference, the deductive ability to identify doublespeak of others, and (2) Deceptive control, the ability to manage one’s information to mislead others through effective doublespeak.
 Second, the non-systematic evaluation issue indicates that previous studies lacked structured methods for error analysis.
@@
 We believe that conducting qualitative analysis in a systematic way provides both reproducibility and interoperability of experimental results.
-Therefore, we adopt a systematic approach known as thematic analysis [11] to identify recurring reasoning errors that hinder LLMs’ doublespeak.
+In this context, we raise our second research question (RQ2): What types of reasoning errors can be systematically identified through qualitative analysis, and how do these errors hinder LLMs from supporting obscured communication?
+To address this question, we employ a systematic approach known as thematic analysis [11] to identify recurring reasoning errors that hinder LLMs’ doublespeak.
 By correlating these qualitative patterns with our quantitative metrics, we uncover how specific reasoning failures impact LLMs’ performance in subtext inference and deceptive control.
-We conducted two studies to address these issues.
 In Study 1 (Section IV), we perform a quantitative analysis using our fine-grained metrics.
@@
 Consequently, this study makes the following contributions:
-  • We demonstrate the effectiveness of using microscopic quantitative analysis to assess the doublespeak of LLMs.
-Specifically, using seven fine-grained metrics, we show possibility of studying doublespeak in LLMs’ SDG play.
-  • We identify four categories of abnormal reasoning patterns that harm obscured communication.
-Also, by combining results from the two studies, we discover a correlation between the results of the two analyses.
+  • We propose a fine-grained evaluation framework for analyzing LLMs’ doublespeak in obscured communication.
+Specifically, our proposed metrics enable the assessment of detailed skills that cannot be evaluated using coarse-grained metrics.
+  • We identify four categories of reasoning failures that hinder LLMs from supporting obscured communication, and show that these categories provide explanatory support for our quantitative results.
 II Related Works
@@
 First, we outline how recent research has shifted its focus from evaluating LLMs on informative communication tasks to more complex settings that require obscured communication.
-Then, we point out two issues posed in their assessment: macroscopic evaluation and non-systematic analysis.
+Then, we point out two limitations posed in their assessment: macroscopic evaluation and non-systematic analysis.
 II-A Adapting LLMs for Obscured Communication
 LLMs have shown human-level performance in various informative communications.
-For example, researchers widely tested LLMs’ ability to question-answering or summarization [12, 13].
+For example, researchers widely tested LLMs’ ability in question-answering or summarization [12, 13].
 To accurately and reliably provide answers, LLMs for informative communication attempt to leverage true knowledge and suppress false statements during the generation process [14, 15].
@@
 Specifically, they adopted particular game conditions, such as win rate or action success rate, that occur during their environmental games as their evaluation metrics.
-However, we suspect that such metrics are too coarse-grained to adequately capture intermediate behaviors during gameplay.
+However, we suspect that such metrics are too coarse-grained to capture intermediate behaviors during gameplay adequately.
 Specifically, the win rate is only available post hoc, limiting its utility for formative assessment.
 Additionally, the action success rate may fail to account for the multi-player dynamics of SDGs in which performance involves not only task execution but also managing others’ perceptions, such as avoiding suspicion.
-In other words, achieving a higher win rate or action success rate in an SDG does not entail that the model possesses superior competence across all sub-skills involved in obscured communication.
+In other words, achieving a higher win rate or action success rate in an SDG does not entail that the model possesses superior competence across all skills involved in core aspects of obscured communication.
 Other researchers proposed more granular metrics, such as logicality or social behaviors, that can be computed by scoring the reasoning or statements made by LLMs during an Avalon game[9, 8].
-However, such metrics proposed in those studies rely on an auxiliary LLM to qualitatively assess reasonings or statements, rather than being grounded in objective in-game events.
+However, such metrics proposed in those studies rely on an auxiliary LLM to qualitatively assess reasoning or statements, rather than being grounded in objective in-game events.
 Therefore, fine-grained quantitative metrics grounded in objective in-game events are essential for evaluating LLMs’ abilities in multi-agent obscured communication, such as subtext inference and deceptive control.
@@
 To enhance the utility of qualitative analysis, it should serve not only to validate the outcomes of quantitative evaluations but also to provide insights into the underlying causes of results.
-In summary, the existing literature has limitations in its quantitative and qualitative analysis methods.
-To discover whether LLMs can support obscured communication, quantitative analysis with fine-grained metrics and validation of qualitative analysis are necessary.
 III Experiment
+TABLE I: Overview of players in SpyGame
+# Player Objectives Required Abilities
+Spy 1 Identify the hidden location while avoiding detection Concealing or Revealing Information
+Citizen 6 Detect the spy without revealing the location
 This section details the process of collecting gameplay logs from LLMs for both quantitative and qualitative analyses.
@@
 III-A1 Team formation
-Similar to Spyfall, SpyGame requires two teams: six citizens and one spy.
+As illustrated in Table I, SpyGame requires two teams: six citizens and one spy.
 The citizens share a common location, called a hidden location, with each citizen playing a specific character in the location.
@@
 From the second turn, the spy player guesses the hidden location by herself (Figure 1c).
-For each turn, she tells her guess to the system secretly and provides a number indicating how certain she is.
+For each turn, she secretly tells the system her guess and provides a number indicating her level of certainty.
 When the certainty is high enough [^2] In our implementation, we use 9 as the threshold out of 10., the system announces her guess and ends the game.
@@
 We selected four LLMs for our experiment: GPT-4[1], GPT-3.5-turbo (ChatGPT), Gemini-pro 1.0[3], and Llama2-70b-chat[2].
-As obscured communication requires reasoning on subtexts, we selected large enough models with sufficient reasoning ability on simple tasks.
+As obscured communication requires reasoning on subtexts, we selected models that are large enough and have sufficient reasoning ability on simple tasks.
 Specifically, we selected models with a parameter size over 70 billion and a reasoning performance exceeding 70% accuracy[21] in 2023.
@@
 Consequently, we collected a total of 168 gameplay logs.
-Table I lists statistics of collected logs.
-On average, each game took 5.4 turns, with 23.8 reasoning steps which refers to a unit of LLM’s reasoning for taking an action.
-TABLE I: Average statistics for all players in the collected logs
+Table II lists statistics of collected logs.
+On average, each game took 5.4 turns, with 23.8 reasoning steps, which refers to a unit of LLM’s reasoning for taking an action.
+TABLE II: Average statistics for all players in the collected logs
 Turns Reasoning Steps Utterances Words/Utterance
@@
 Total 5.45 23.86 13.31 23.07
-IV Microscopic Quantitative Analysis
-Our study aimed to investigate whether LLMs can support obscured communication by conducting a microscopic analysis of their SDG playing.
-We designed seven metrics that enable fine-grained analysis of two key SDG facets: subtext inference and deceptive control.
-Compared with most metrics adopted by previous research, which were designed on game-level outcomes, our approach allows for an analysis of intermediate behaviors during the game.
-Also, our quantitative metrics are fully automated metrics that do not require any manual labeling.
-In this section, we first discuss whether existing simple metrics are sufficient for measuring intermediate behaviors of LLMs.
-Then, we demonstrate our analysis can discover new insights that could not be captured by using current metrics by discussing our quantitative results.
-IV-A Insufficiency of Previous Analysis
+IV Quantitative Analysis with Fine-Grained Metrics
+TABLE III: Overview of tested metrics
+Metric Precondition Level Measuring Point
+Prior Win Rate - Coarse-grained Outcome of game
+Metrics Living Round - Coarse-grained How long survived
+Guess Success Rate Spy win Mid-grained How subtext inference ability contributed to win
+Information Catching Rate Exposure occured Fine-grained How successfully exploited explicit clue
+Our Information Deduction Rate Exposure not occured Fine-grained How successfully exploited implicit clue
+Metrics Caught Rate Spy lose Mid-grained How deceptive control ability contributed to lose
+Vote Rate Day/Final vote occured Fine-grained How successfully avoided concentrated suspicion
+Vote Entropy Day/Final vote occured Fine-grained How successfully dispersed suspicion
+To investigate RQ1 regarding the adequacy of coarse-grained metrics in capturing LLMs’ abilities, we conduct a fine-grained analysis of model behavior in SDG.
+To this end, we design six quantitative metrics, which are presented in Table III, targeting two core aspects of obscured communication subtext inference and deceptive control.
+For each aspect, we propose one mid-grained metric to assess its overall contribution to the spy’s success or failure, and two fine-grained metrics to evaluate how well the model performs the associated skills.
+From the following subsection, we first discuss the limitations of existing simple metrics, which focus solely on game-level outcomes and are insufficient for capturing the intermediate behaviors exhibited by LLMs during gameplay.
+We then demonstrate how our proposed metrics uncover insights into the way LLMs utilize specific reasoning skills in obscured communication, based on our quantitative analysis.
+Additionally, we performed statistical analyses to assess the degree of model-level differences, with the expectation that such variation may offer additional interpretive insights.
+To examine differences between models within each metric, we conducted statistical analyses.
+Given the small size of 168 games, we used non-parametric statistical tests.
+For metrics composed of continuous variables, we performed the Kruskal-Wallis test[23] to examine group differences.
+For metrics consisting of categorical variables, we used the chi-squared test[24] to assess differences between groups.
+IV-A Limitations of Existing Metrics
 Previous studies commonly adopted macroscopic analyses to examine SDG gameplay behavior[8, 20, 19, 7, 6].
-The most frequently used metrics are the win rates (WR) and living rounds (LR).
+The most frequently used metrics are the win rate (WR) and living round (LR).
 These metrics evaluate a player’s performance based on final game outcomes: WR is determined by the win/loss result, while LR reflects how long the player remained active during the game.
   • Win Rate (WR) is the percentage of games won by the spy out of the total number of games.
+    $WR=\frac{\#\text{ victory}}{\#\text{ total games}}$ (1)
   • Living Round (LR) is the average number of turns the spy survived, across all games.
+    $LR=\frac{\sum\#\text{ turns spy survived}}{\#\text{ total games}}$ (2)
 IV-A1 Metric result
-Tables II (a) and III (a) present the results of the coarse-grained metrics when playing against strong and weak citizens, respectively.
-The WR and LR exhibited different trends in SDG play performance.
-In terms of WR, GPT-4 showed the highest performance against strong citizens (80.95%), followed by Gemini (57.14%), Llama (52.38%), and GPT-3.5 (33.33%).
-A similar order of performance was observed for against weak citizens.
-However, in terms of LR, Llama demonstrated the highest performance against strong citizens (9.00), followed by Gemini (6.95), GPT-4 (5.80), and GPT-3.5 (3.71).
-This trend is similar when examining the performance playing against weak citizens.
-Note that these metrics yield different results regarding the best-performing model: WR indicates GPT-4 as the best, while LR identifies Llama.
-This discrepancy is due to multiple winning conditions specified in the SDG.
-These conditions allow players to win games in several ways.
-Therefore, to accurately identify the skills that contribute to winning, it is necessary to consider multiple abilities simultaneously.
-However, since WR and LR consider only the final state of the game, these metrics cannot examine different abilities separately.
-For example, in SpyGame, LR indicates how well a spy survives during the game.
-However, a longer survival does not always guarantee victory, leading to potential discrepancies between LR and WR.
+Tables V (a) and VI (a) present the results of the coarse-grained metrics when LLMs played against strong and weak citizens, respectively.
+When playing against strong citizens, statistical differences across models were found in both WR ($\chi^{2}=9.81$, $p<0.05$) and LR ($H=42.09$, $p<0.001$) metrics: GPT-4 achieved the highest WR (80.95%), followed by Gemini (57.14%), Llama (52.38%), and GPT-3.5 (33.33%).
+However, when playing against weak citizens, no significant difference was found in WR across models ($\chi^{2}=4.14$, $p=0.24$).
+We suspect that weak citizens’ low gameplay proficiency made it easier for all models to achieve high win rates as the spy.
+Unlike WR, LR revealed a different trend.
+Llama consistently recorded the longest survival time (9.00 and 7.85), even outperforming GPT-4, which had the highest WR.
+This discrepancy between WR and LR illustrates that success in surviving longer does not necessarily translate to overall victory.
+It suggests that each metric captures different aspects of gameplay, potentially tied to distinct skills
+Furthermore, this discrepancy highlights that coarse-grained metrics are challenging to interpret because both WR and LR are influenced by a range of overlapping factors, including guessing accuracy, voting dynamics, and the amount of information citizens disclose.
+As a result, it remains unclear which specific underlying skills these metrics are actually measuring.
 Hence, a microscopic method is required for a comprehensive analysis of LLM’s performance in various facets.
-TABLE II: Quantitative results for playing against strong citizens
-(a) General (b) Subtext inference (c) Deceptive Control
-Winning Living Guess Notice Information Information Caught Vote Vote
-Model Rate (%) Round (#) Success (%) Rate (%) Catching Deduction (%) Rate (%) Rate (%) Entropy
-Gemini 57.14 6.95 50.00 $3$ 30.00 022.22 35.7 0.348
-(14) (20) (9) (42)
-GPT-3.5 33.33 3.71 27.78 28.57 - 28.57 007.14 25.0 0.730
-(18) (21) (14) (12)
-GPT-4 80.95 5.80 81.25 61.90 - 61.90 025.00 33.3 0.953
-(16) (21) (4) (30)
-Llama2 52.38 9.00 - 09.52 - 00.00 100.00 45.2 0.634
-(0) (19) (10) (126)
-Total 55.95 6.365 35.78 $3$ 30.86 037.83 40.47 0.642
-Metrics whose lower values indicate better results.
-Note that Information Catching was not measured playing against strong citizens.
-TABLE III: Quantitative results for playing against weak citizens
-(a) General (b) Subtext inference (c) Deceptive control
-Winning Living Guess Notice Information Information Caught Vote Vote
-Model Rate (%) Round (#) Success (%) Rate (%) Catching (%) Deduction Rate (%) Rate (%) Entropy
-Gemini 90.48 5.28 85.71 090.48 61.11 - 00.00 07.1 0.420
-(14) (18) (2) (42)
-GPT-3.5 71.43 2.52 71.43 076.19 92.86 - $0$
-(21) (14) (6) (0)
-GPT-4 90.48 2.42 90.48 100.00 87.50 - $0$
-(21) (16) (2) (0)
-Llama2 76.19 7.85 60.00 071.43 20.00 - 60.00 13.5 0.605
-(5) (15) (5) (96)
-Total 82.14 4.51 80.32 084.52 65.07 - 20.00 11.59 0.548
-Metrics whose lower values indicate better results.
-Note that Information Deduction was not measured playing against weak citizens.
-IV-B Subtext Inference
-We designed multiple metrics to enable a microscopic analysis of subtext inference.
-We operationally define subtext inference as the ability to uncover hidden information through deduction or by capturing pieces of information from the conversation.
-Subtext inference has two sub-abilities: information capture and information deduction.
-Information capture measures whether a spy notices and exploits critical hints inadvertently revealed by citizens.
-Since weak citizens generally leak such information unintentionally, information capture ability becomes essential for the spy to win the game against them.
-Since strong citizens rarely make such mistakes, it is highly challenging to observe information capture of a spy.
-On the other hand, information deduction mainly focuses on how effectively a spy infers a hidden location by combining various pieces of information.
-Since strong citizens generally conceal the key information, information deduction ability becomes essential for the spy to uncover information embodied in a citizen’s utterance.
+IV-B Analysis of Subtext Inference
+We designed three metrics to enable a microscopic analysis of the first core aspect, subtext inference.
+This aspect concerns the model’s ability to uncover hidden information, either through deduction or by capturing clues embedded in the conversation.
+In particular, subtext inference has two skills: information capture and information deduction.
+Information capture measures whether a spy notices and exploits explicit clues inadvertently revealed by citizens.
+Since weak citizens generally leak such information unintentionally, the spy’s ability to capture information becomes essential to win the game against them.
+On the other hand, information deduction mainly focuses on how effectively a spy infers the hidden location by synthesizing implicit clues that citizens intentionally crafted to signal their identity to other citizens without revealing the location.
+Since strong citizens generally conceal key information, the skill of deducing information becomes essential for the spy to uncover the information embodied in a citizen’s utterance.
 To evaluate both information capture and deduction, we analyzed the outcomes of the spy’s guessing.
@@
 IV-B1 Metric design
-To assess the spy’s ability to identify hidden intent, we propose four metrics based on their guessing behavior. These metrics range from relatively summative to more formative.
-  • Guess Success (GS) measures how much the guessing of a spy contributed to their victory.
+To evaluate how subtext inference contributes to the spy’s success and to diagnose the state of its two underlying skills, we propose three metrics.
+TABLE IV: Multifaceted interpretation of subtext inference
+Information Information Case Interpretation
+Catching Deduction
+Ideal High High The model effectively utilizes both sub-skills and demonstrates ideal subtext inference
+Moderate High Low The model captures implicit clues well but struggles to synthesize them into a final deduction
+Moderate Low High The model shows strong reasoning ability but fails to identify explicit clues from the conversation
+Worst Low Low The model fails to capture and reason over subtext, indicating poor subtext inference ability
+  • Guess Success Rate (GSR) measures how much the guessing of a spy contributed to their victory.
 This metric focuses only on games where the spy attempted to guess, rather than all played games.
-We calculate GS by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted.
-A higher GS indicates that the spy has successfully identified hidden locations, demonstrating strong subtext inference skills.
-  • Notice Rate (NR) indicates how well a spy noticed the hidden location across all games.
-It is calculated by computing the percentage of games in which the spy correctly detected the hidden location over the total number of games played.
-Unlike GS, which is measured only when the spy decides to reveal their identity publicly, NR can be computed across all games.
-This is because NR mainly considers whether the spy notices the hidden location, regardless of whether they decide to guess.
-  • Information Catching (IC) measures how often the spy successfully leverages exposed information to achieve a victory.
-Unlike NR, which only assesses the ability of the spy to identify the hidden location, IC evaluates the effectiveness in employing such information to win.
-We specifically focus on games where weak citizens directly exposed information about the hidden location instead of considering all the played games.
-Here, the exposure means a case where a citizen’s utterance literally contains the hidden location [^5] Qualitative analysis in Section V-B1 discusses the case..
-We measured IC by computing the proportion of games in which the spy successfully guessed the hidden location from games where exposure occurred.
-Thus, a high IC score reflects the ability of the spy to achieve victory by effectively capturing and utilizing key information.
-  • Information Deduction (ID) is similar to IC but it applies to games where exposure did not occur.
-This metric measures how well the spy combined pieces of information to determine the hidden location.
-To compute ID, we divided the number of games in which the spy successfully guessed the hidden location by the number of games without any exposure.
-Therefore, a high ID indicates that the spy successfully combined clues from the conversation and utilized them to achieve victory.
-We provide a multifaceted interpretation of subtext inference by jointly analyzing IC and ID metrics.
+We calculate GSR by dividing the number of games in which the spy made a successful guess by the number of games in which guessing was attempted.
+A higher GSR suggests that the spy is more capable of inferring subtextual information to reach out to hidden targets.
+    $GSR=\frac{\#\text{ victory by }Guessing}{\#\text{ victory}}$ (3)
+  • Information Catching Rate (ICR) measures how effectively the spy exploits explicit clues directly stated by citizens to identify the hidden location.
+This metric focuses only on games in which exposure occurred rather than all games.
+We specifically focus on games where exposure occurred rather than all the played games.
+Here, the exposure refers to a case where a citizen’s utterance literally contains the hidden location [^5] Qualitative analysis in Section V-B1 discusses the case..
+[…]
+  • Information Deduction Rate (IDR) measures whether the spy can exploit implicit clues, which are scattered hints embedded across multiple citizen utterances, to infer the hidden location without exposure.
+This metric applies only to games where exposure did not occur, meaning the hidden location was never explicitly mentioned by any citizen.
+Unlike ICR, where success may depend on detecting an obvious clue, IDR reflects the spy’s capacity to synthesize indirect linguistic signals and reason under ambiguity.
+We compute IDR as the proportion of games without exposure in which the spy correctly guessed the location.
+[…]
+As illustrated in Table IV, we provide a multifaceted interpretation of subtext inference by jointly analyzing ICR and IDR metrics.
 Based on whether each metric is above or below its average, we classify model behavior into three combinations.
-The first combination, when both IC and ID are high, represents an ideal scenario in which the spy successfully identifies the hidden location regardless of the type of citizen.
-The second combination occurs when only one of the metrics, either IC or ID, is higher than the average.
-This suggests a relative lack of proficiency in the corresponding sub-ability.
-For instance, a model with high IC but low ID demonstrates strong information capture but weak information deduction.
-The final and least favorable combination for subtext inference is when both IC and ID are low.
+The first combination, when both ICR and IDR are high, represents an ideal scenario in which the spy successfully identifies the hidden location regardless of the type of citizen.
+The second combination occurs when only one of the metrics, either ICR or IDR, exceeds the average.
+This suggests a relative lack of proficiency in the corresponding skills.
+For instance, a model with high ICR but low IDR demonstrates strong information capture but weak information deduction.
+The final and least favorable combination for subtext inference is when both ICR and IDR are low.
 This combination indicates that the spy failed to infer the hidden location in most games.
-Moreover, by combining GS and NR, we can explore whether the spy effectively incorporates the incoming information to maintain their guess.
-The ideal combination is when both GS and NR are high.
-This indicates that the spy consistently maintained a correct guess after identifying the hidden location, until the end of the game.
-In contrast, a high GS paired with a low NR suggests that while the spy struggled to predict the hidden location initially, they managed to maintain their guess once the correct location was identified.
-Meanwhile, the combination of low GS and high NR implies that although the spy successfully guessed the hidden location, the spy is prone to discarding the guess easily.
-The worst-case scenario is when both NR and GS are low, indicating a significant failure in both identifying and maintaining the correct hidden location.
+This categorization allows us to diagnose which skill each model lacks, providing insight into how the model could be improved to more effectively support obscured communication.
 IV-B2 Result
-Tables II (b) and III (b) present the measurements for subtext inference when playing against strong and weak citizens, respectively.
-The results indicate that GPT-4 has sufficient ability compared to the other models.
-GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (81.25, 50.0, 27.78, and 0%, respectively).
-The results in terms of NR show a similar trend: GPT-4 (61.90%), Gemini (33.33%), GPT-3.5 (28.57%), and Llama (9.52%).
-Finally, regarding ID, GPT-4 was the best model, followed by Gemini, GPT-3.5, and Llama (61.90, 30.0, 28.57, and 0%).
-In summary, GPT-4 demonstrated the best performance against strong citizens in subtext inference.
-Meanwhile, Llama showed the lowest performance against strong citizens.
-Against weak citizens, GPT-4 was the best model in terms of GS, followed by Gemini, GPT-3.5, and Llama (90.48, 85.71, 71.43, and 60%, respectively).
-NR displayed the same trend: GPT-4 (100%), Gemini (90.48%), GPT-3.5 (76.19%), and Llama (71.43%).
-Lastly, regarding IC, GPT-3.5 (92.86%) showed the best performance, differing in GS and NR, followed by GPT-4 (87.5%), Gemini (61.11%), and Llama (20.0%).
-Similar to the results against strong citizens, GPT-4 showcased good performance against weak citizens.
+Tables V (b) and VI (b) present the measurements for subtext inference when playing against strong and weak citizens, respectively.
+When playing against strong citizens, we found there were statistically significant differences between LLMs in all metrics regarding the subtext inference, including Guess Success Rate (GSR, $\chi^{2}=8.68$, $p<0.05$) and Information Deduction Rate (IDR, $\chi^{2}=18.02$, $p<0.001$).
+Specifically, GPT-4 was the best model in terms of GSR, followed by Gemini, GPT-3.5, and Llama (81.25, 50.00, 31.57, and 0%).
+This result indicates that GPT-4 has sufficient ability compared to other models in achieving victory by utilizing collected information during the games.
+Similarly, GPT-4 also exhibited the best score in IDR, followed by Gemini, GPT-3.5, and Llama (61.90, 30.0, 28.57, and 0%).
+This suggests that GPT-4 possesses a strong ability to leverage implicit clues to reach out to the hidden location.
+Against weak citizens, while the overall ranking trend among models for the Guess Success Rate (GSR) remained consistent, no statistically significant differences were observed.
+We attribute this to the fact that weak citizens frequently revealed explicit clues that allowed the spy to easily infer the hidden location.
+In contrast, we found statistically significant differences between the tested models in Information Catching Rate (ICR, $\chi^{2}=21.83$, $p<0.001$).
+Specifically, GPT-3.5 outperformed other models (92.86%), followed by GPT-4 (87.50%), Gemini (61.11%), and Llama (20.00%).
+This suggests that GPT-3.5 is particularly effective at exploiting explicit clues often leaked by weak citizens.
 IV-B3 Discussion
+TABLE V: Quantitative results for playing against strong citizens
+(a) Prior Metric (b) Subtext inference (c) Deceptive Control
+Winning Living Guess Information Information Caught Vote Vote
+Model Rate Round Success Rate Catching Rate Deduction Rate Rate Rate Entropy
+Gemini 57.14 6.95 50.00 - 30.00 022.22 35.71 0.348
+(14) (20) (9) (42)
+GPT-3.5 33.33 3.71 31.57 - 28.57 007.14 25.00 0.730
+(19) (21) (14) (12)
+GPT-4 80.95 5.80 81.25 - 61.90 025.00 33.33 0.953
+(16) (21) (4) (30)
+Llama2 52.38 9.00 - - 00.00 100.00 46.03 0.657
+(0) (19) (10) (126)
+Total 55.95 6.36 53.06 - 30.86 037.83 40.95 0.642
+Metrics whose lower values indicate better results.
+Note that Information Catching cannot be   measured playing against strong citizens.
+TABLE VI: Quantitative results for playing against weak citizens
+(a) Prior Metric (b) Subtext inference (c) Deceptive control
+Win Living Guess Information Information Caught Vote Vote
+Model Rate Round Success Rate Catching Rate Deduction Rate Rate Rate Entropy
+Gemini 90.47 5.28 85.71 61.11 - 00.00 07.14 0.420
+(14) (18) (2) (42)
+GPT-3.5 71.42 2.52 71.42 92.86 - 00.00 - -
+(21) (14) (6) (0)
+GPT-4 90.47 2.42 90.47 87.50 - 00.00 - -
+(21) (16) (2) (0)
+Llama2 76.19 7.85 60.00 20.00 - 60.00 15.62 0.605
+(5) (15) (5) (96)
+Total 82.14 4.52 80.32 65.07 - 20.00 13.04 0.548
+Metrics whose lower values indicate better results.
+Note that Information Deduction was cannot be   measured playing against weak citizens.
 In this section, we derive insights using the proposed metrics by conducting a multifaceted analysis.
-This discussion encompasses three aspects: (1) a complementary relationship between IC and ID; (2) an exploration of the spy’s filtering behavior on incoming information; and (3) the need for formative metrics concerning subtext inference.
-First, IC and ID are complementary metrics that enable multifaceted analysis of subtext inference.
-As aforementioned, these two metrics can be combined to form three distinct partitions, each representing different levels of subtext inference ability.
-Using these partitions, we can determine the best model possessing sufficient sub-abilities for subtext inference; In our analysis, GPT-4 demonstrated high scores in both IC and ID, with its ID being double that of the second-place model.
-Meanwhile, GPT-3.5 falls within the high-IC region but not the high-ID region.
-This indicates that while GPT-3.5 struggles to combine hints from the conversation to deduce the hidden location.
-Second, by analyzing GS and NR, we can examine how the spy filters incoming information to maintain their guess.
-As suggested in the metric design, we can use four quadrants made by GS and NR.
-Among the four models, GPT-4 belongs to the first partition and achieves a high score on both metrics.
-This indicates that GPT-4 can filter useful information, reinforcing the belief of the spy in the hidden location.
-This could be a hint that GPT-4 has strong information deduction sub-ability.
-Meanwhile, belonging to different partitions can induce differences in such filtering behaviors.
-For example, although Gemini and GPT-3.5 exhibited similar NR scores against strong citizens, GS of Gemini was almost double that of GPT-3.5.
-This difference suggests that Gemini is more selective in using guess actions than GPT-3.5, as evidenced by the fact that Gemini used 14 guess actions regardless of citizens, whereas GPT-3.5 utilized 21 (against weak) and 18 (against strong) guess actions.
+This discussion encompasses three aspects: (1) a complementary relationship between Information Catching Rate (ICR) and Information Deduction Rate (IDR); (2) revealing decision-making behaviors through comparative analysis of metric-level disparities; and (3) the need for formative metrics concerning subtext inference.
+First, ICR and IDR are complementary metrics that enable multifaceted analysis of subtext inference.
+By using the categorizations illustrated in Table IV, we can determine the best model possessing sufficient skills for subtext inference; In our analysis, GPT-4 exhibits strong performance in both ICR and IDR, with its IDR score being twice as high as that of the second-best model.
+In contrast, GPT-3.5 shows high ICR but low IDR, suggesting that while it is effective at capturing explicit clues, it struggles to infer the hidden location when only subtle, implicit cues are available.
+This observation suggests that for models exhibiting a similar profile–high ICR but low IDR–future efforts to improve their ability to support obscured communication should focus on enhancing their capacity to infer hidden meanings from subtle contextual cues, rather than relying on explicit exposures.
+Second, our metrics reveal a discrepancy between LLMs’ ability to collect information and their decision-making thresholds for acting on the information.
+While tested models differ significantly in their ability to gather information (ICR and IDR), their thresholds for acting appear more consistent.
+Specifically, although the maximum gaps in GSR across models are 49.68 and 30.47% against weak and strong citizens, respectively, the corresponding gaps in ICR and IDR are much larger: 72.86 and 61.90%, respectively.
+In addition to these numeric differences, statistical tests further confirm that model differences in ICR and IDR are more pronounced than in GSR.
+This suggests that although LLMs vary greatly in how well they capture or synthesize clues from context, they tend to behave similarly when deciding whether enough evidence has been gathered to initiate a guess.
+In other words, the threshold for taking action based on inferred information appears more uniformly developed than the underlying abilities to process that information.
 Finally, our metrics regarding subtext inference enable more detailed analyses of the factors contributing to the scores of simple metrics.
@@
 Against strong citizens, Gemini showed a similar WR to Llama.
-Meanwhile, against weak citizens, Gemini exhibited the same WR as GPT-4.
-However, these observations are somewhat shallow when we consider our metrics, as the three models display different subtext inference abilities.
-For example, against strong citizens, Gemini had a higher ID score than Llama in that Llama completely failed to combine clues about the hidden location.
-Similarly, against weak citizens, we can conclude that GPT-4 is slightly better than Gemini because of its higher GS and NR scores.
+Meanwhile, against weak citizens, Gemini exhibited the same numerical WR score as GPT-4.
+However, these observations are somewhat shallow when considering our metrics, as the three models display different subtext inference abilities.
+For example, against strong citizens, Gemini had a higher IDR score than Llama, as Llama completely failed to synthesize implicit clues about the hidden location.
+Similarly, against weak citizens, we can conclude that GPT-4 is better than Gemini because of its higher ICR score.
 Thus, proposed metrics can reveal finer differences between LLMs than existing metrics.
-In addition, such a detailed analysis revealed that GPT-3.5 is overly confident in its deduction results.
-As illustrated in Tables II and III, the spy with GPT-3.5 demonstrated relatively lower NR scores; however, the spy used guessing actions more frequently.
-That is, the spy rushed into making guesses even when their inference about the hidden location was incorrect.
-One possible explanation is that GPT-3.5 may treat its own inference as hard evidence, as we will discuss in Section V.
-We suspect that this mechanism may have contributed to GPT-3.5 achieving the lowest WR score, despite its relatively good subtext inference ability.
-IV-C Deceptive Control
-In addition to subtext inference, we designed four metrics for the multifaceted analysis of deceptive control.
-Here, we operationally define deceptive control as the ability to blend in with others by controlling information.
-Deceptive control encompasses two sub-abilities: evading suspicion and dispersing suspicion.
-Evading suspicion mainly considers how well the spy avoids being suspected by others.
+IV-C Analysis of Deceptive Control
+In addition to subtext inference, we designed three metrics for the second core aspect, deceptive control.
+This aspect concerns the model’s ability to blend in with others by controlling information.
+Deceptive control encompasses two skills: evading suspicion and dispersing suspicion.
+Evading suspicion mainly involves how well the spy avoids being suspected by others.
 It is quantified by the number of suspicions directed at the spy.
@@
 The following subsections describe the design of the four metrics and their results.
-The results indicate that GPT-3.5 is the most powerful spy player, followed by GPT-4, Gemini, and Llama.
 IV-C1 Metric design
 Based on day/final votes and accusations during a game, we suggest three metrics for measuring deceptive control, ranging from relatively summative to highly formative.
-  • Caught Rate (CR) is how frequently the spy is caught by citizens in the Day/Final vote.
-We compute this by dividing the number of games in which the spy was caught by the number of games lost.
+  • Caught Rate (CR) measures how much the voting of citizens contributed to the spy’s failure.
+We compute this by dividing the number of games in which the spy was caught by the number of games the spy lost.
 Thus, a high CR indicates that the majority of citizens easily identified the spy player during the game.
 Note that this is a summative metric for the spy’s deceptive control ability because it only considers the final state of the game.
-  • Vote Rate (VR) is the proportion of suspicions about the spy across all votes that occurred.
-We measure this by computing the percentage of votes the spy received from citizens across the total number of votes, excluding the spy’s votes.
+    $CR=\frac{\#\text{ defeat by }Voting}{\#\text{ defeat}}$ (6)
+  • Vote Rate (VR) measures how much suspicion was concentrated on the spy during a game.
+We calculate it as the proportion of total votes that citizens cast against the spy in each game.
 Therefore, a high VR indicates that the spy failed to evade suspicion during the game.
+[…]
   • Vote Entropy (VE) is how scattered the suspicions about the spy are in the game.
-We measure this by computing the entropy of the votes that citizens casted for each game and averaging them.
+We measure this by computing the entropy of the votes that citizens cast for each game and averaging them.
 Note that in SpyGame, a player’s suspicion cannot affect another player’s suspicion because they cannot have a discussion together.
@@
 Thus, the value of VE increases.
-We devised a multifaceted interpretation of deceptive control by using VR and VE simultaneously.
-By partitioning VR and VE by their average values, we derived four combinations.
-First, the most balanced combination was a low VR and high VE.
-This is an ideal combination because the spy can successfully evade others’ suspicions and lead citizens to discourage each other from narrowing down spy candidates.
-Second, models in which VR and VE are both low or high lack one of the two sub-abilities.
-Finally, the worst combination occurs when VR is high and VE is low.
-This combination indicates that citizens’ suspicions were concentrated on the spy and that the spy failed to undermine suspicions.
+    $VE=\frac{\sum(Entropy(\text{votes cast by citizens))}}{\#\text{games with }Voting}$ (8)
+As illustrated in Table VII, we developed a multifaceted interpretation of deceptive control by jointly analyzing Vote Rate (VR) and Vote Entropy (VE).
+By partitioning both metrics based on their average values, we identified four representative combinations.
+First, the most desirable combination is low VR and high VE, which reflects an ideal form of deceptive control: the spy avoids being directly targeted while simultaneously inducing citizens to spread their suspicions across each other.
+Second, combinations where both VR and VE are either low or high suggest an imbalance or partial deficiency in the skills involved in deceptive control.
+For example, high values in both metrics indicate that although suspicion is spread among citizens, the spy remains a frequent target, which reflects only partial success in deception.
+Finally, the worst-case scenario is high VR and low VE, where the spy becomes the focal point of suspicion and fails to redirect or confuse citizens.
 IV-C2 Result
-Tables II (c) and III (c) elaborate measurements of deceptive control when playing against strong and weak citizens, respectively.
-Note that some missing values for VR and VE indicate that the spy guessed the hidden location for all games, and no votes were conducted.
-The results indicate that the variants of GPT have sufficient ability compared to other models.
-Against strong citizens, GPT-3.5 was the best model in terms of CR, followed by Gemini, GPT-4, and Llama (7.1, 22.2, 25.0, and 100%, respectively).
-The results for VR showed a similar tendency: GPT-3.5 (25%), GPT-4 (33.3%), Gemini (35.7%), and Llama (45.2%).
-Finally, regarding VE, GPT-4 was the best model, followed by GPT-3.5, Llama, and Gemini (0.953, 0.730, 0.634, and 0.348, respectively).
-Interestingly, GPT-3.5 showed deceptive control performance similar to that of GPT-4, although GPT-4 showed much higher performance in other language-based tasks[1].
-Against weak citizens, GPT variants showed good performance compared to others.
-Because GPT variants ended the game by guessing, citizens had no chance to find the spy.
-Thus, CR values for GPT-3.5 and GPT-4 were zero, and VR and VE could not be measured using these models.
-Other models showed a tendency similar to that against strong citizens.
-Regarding CR and VR, Gemini (0% and 7.1%, respectively) was better than Llama (60% and 13.5%).
-However, regarding VE, Gemini (0.420) performed worse than Llama (0.605), as shown by results against strong citizens.
+TABLE VII: Multifaceted interpretation of deceptive control
+Vote Vote Case Interpretation
+Rate Entropy
+Ideal Low High The model successfully avoids direct suspicion while distributing suspicion among citizens
+Moderate High High Although the model encourages suspicion to be spread, it still attracts frequent votes
+Moderate Low Low The model avoids being voted, but fails to diffuse suspicion among citizens
+Worst High Low The model becomes the primary target of suspicion and fails to spread suspicion across other players
+Tables V (c) and VI (c) elaborate measurements of deceptive control when playing against strong and weak citizens, respectively.
+Note that some missing values for Vote Rate (VR) and Vote Entropy (VE) indicate that the spy guessed the hidden location for all games, and no votes were conducted.
+Against playing strong citizens, we found statistically significant differences between LLMs in Caught Rate (CR, $\chi^{2}=23.25$, $p<0.001$) while no differences in VR or VE.
+Specifically, GPT-3.5 achieved the lowest CR (7.14%), followed by Gemini, GPT-4, and Llama (22.22, 25.00, and 100.00%).
+Since CR measures how much the voting of citizens contributed to the spy’s failure, this result suggests that GPT-3.5 was rarely identified as the spy when it lost.
+Meanwhile, VR and VE did not show statistically significant differences across models, although the overall ranking trends were similar to those of CR.
+We interpret the small differences in VR and VE as a reflection of how models with high CR, such as Llama, were often caught after receiving a minimal majority of votes in every voting.
+In other words, although they lost, they were not overwhelmingly targeted against strong citizens.
+Against weak citizens, no statistically significant differences were observed across models in any of the deceptive control metrics.
+We attribute the lack of difference in CR to the overall high winning rates, which resulted in very few lost games and thus limited opportunities for the spy to be caught via votes.
+Due to the early termination of most games, voting occurred infrequently.
+As a result, VR and VE could not be measured for GPT variants.
+These observations suggest that, when playing against weak citizens, the spy’s success depends less on the ability to deceive opponents and more on the ability to exploit explicit clues to directly guess the hidden location.
 IV-C3 Discussion
-In this section, we discuss how to derive insights using the proposed metrics to conduct a multifaceted analysis of deceptive control.
-This discussion includes the following two aspects: (1) a complementary relationship between Vote Rate (VR) and Vote Entropy (VE), and (2) additional perspectives captured by adopting formative metrics.
-First, VR and VE are complementary metrics that enable a multifaceted analysis of deceptive control.
-As suggested above, these two metrics can form three partitions explaining the different levels of capacity regarding deceptive control.
-Using these partitions, we can determine the best model for balancing two sub-abilities of deceptive control; GPT family showed gameplay balanced between VR and VE.
-These models can evade and disperse suspicion to win the game.
-In contrast, the worst model in terms of deceptive control was Llama.
-Although Llama performed better against weak citizens than strong citizens, Llama player was often suspected as a spy and disregarded the fact that the majority of players doubted him/her.
-So, the two sub-abilities are both good or both bad in these two cases.
-Gemini showed a different behavior: the model was good at evading suspicion but struggled to disperse suspicion.
-Thus, as illustrated above, VR and VE can provide a wider picture of deceptive control behavior.
-Second, proposed metrics provide additional perspectives on investigating deceptive control, compared to simple summative metrics.
-In section IV-A, we observed a discrepancy between two summative metrics, win rate (WR) and living round (LR).
-For example, both GPT-3.5 and Llama showed relatively low WR scores, but regarding LR, they achieved relatively low and high performance, respectively.
-So, using summative metrics only, we may conclude that Llama outperforms GPT-3.5 in SpyGame.
-However, the proposed metrics suggested different results; GPT-3.5 is good at controlling information, but Llama is not.
-This example shows that our proposed metrics are more suitable for examining deceptive control ability than existing summative metrics.
-Similarly, our proposed metrics are helpful in identifying gameplay problems.
-For example, both GPT-4 and GPT-3.5 showed relatively low LR scores; however, WR achieved the best and worst scores, respectively.
-Therefore, using only summative metrics, we can conclude that GPT-4 outperforms GPT-3.5 in SpyGame.
-However, the proposed metrics suggest different results: both GPT-4 and GPT-3.5 are good at deceptive control regarding VR and VE.
-V Qualitative Analysis
-To reconfirm the findings of quantitative analysis from another perspective, we conducted a thematic analysis.
-We defined three types of reasoning errors in LLM reasoning while deciding their actions by coding abnormal patterns of 15 randomly selected gameplay logs.
+As aforementioned, we found no model achieved a statistically significant advantage, except for Guess Success (GS) when against strong citizens.
+Nevertheless, several cautious insights emerge from the patterns observed in experimental results.
+In this section, we outline two aspects of these findings: (1) the additional perspective gained from fine-grained deception metrics versus coarse-grained metrics, and (2) differences in how LLMs manage suspicion.
+First, the formative metrics provide a complementary perspective beyond outcome-level summative metrics.
+Specifically, evaluating based solely on the coarse-grained metrics, illustrated in V(a) and VI(a) (WR and LR), may mislead conclusions about an LLM’s ability to support a core aspect of obscured communication.
+For instance, an evaluation based solely on WR and LR may misleadingly suggest that Llama2 demonstrates superior performance to GPT-3.5 across experiments.
+In contrast, our fine-grained analysis offers a more nuanced interpretation: Llama2 recorded exceptionally high Caught Rate (CR) scores of 100% and 60% against strong and weak citizens, respectively.
+These findings reflect a deficiency in Llama’s capacity to support deceptive control, as the model was frequently subjected to decisive levels of suspicion, resulting in direct defeat.
+A similar misinterpretation arises when comparing Gemini and Llama using coarse-grained metrics.
+Although both models achieved comparable WR, Llama exhibited a much higher average LR, which might misleadingly imply that the model succeeded in deceiving citizens and thereby extended its survival.
+However, our metrics show that Llama also had a substantially higher CR, meaning it was frequently the primary target of suspicion.
+This implies that more prolonged survival may not necessarily reflect successful deception, but rather passive or non-committal behavior that delays elimination without effectively evading suspicion.
+Such cases highlight a potential risk in using metrics like living round as reward signals in RL-based optimization, as they could incentivize specification gaming by encouraging models to optimize for the metric without actually demonstrating the intended deceptive behavior[25].
+Second, the multifaceted interpretation framework presented in Table VII enables a multifaceted interpretation of how models manage suspicion during gameplay.
+Although the differences are not statistically significant, each model exhibits a distinct pattern across these two dimensions.
+GPT-4 and GPT-3.5 demonstrate relatively low VR and high VE, aligning with the ideal pattern where the spy both avoids being directly targeted and induces dispersed suspicion.
+In contrast, Llama shows a higher VR and lower VE, indicating concentrated suspicion and a lack of effective misdirection.
+Gemini is assigned to the moderate case, maintaining a low VR but struggling to distribute suspicion broadly.
+These results suggest that while coarse metrics obscure such differences, formative metrics can reveal the specific ways in which each model either succeeds or struggles to support deceptive control.
+V Qualitative Analysis via Thematic Coding
+To investigate RQ2 regarding the reasoning errors that hinder LLMs, we conducted a thematic analysis.
+We defined three types of reasoning errors in LLM reasoning when deciding their actions by coding abnormal patterns in 15 randomly selected gameplay logs.
 As a result, we found that these patterns are related to psychological concepts, including identity confusion, memory distortion, and dissociation[26].
-In addition, we found that these concepts are related to the results of the quantitative analysis, especially for caught rate (CR) and guess success (GS).
-The following subsections illustrate the procedure and results of this qualitative analysis.
-V-A Analysis Procedure
+Furthermore, these reasoning errors appear to explain concrete failure modes observed in our quantitative metrics, particularly in Caught Rate (CR) and Guess Success Rate (GSR).
+This alignment between qualitative and quantitative findings helps answer RQ2 by showing not only what types of reasoning errors occur, but also how they directly undermine the LLM’s ability to avoid detection or correctly infer hidden information.
+V-A Qualitative Coding Procedure
 Through thematic analysis, we labeled the abnormal patterns of players in the gameplay logs.
@@
 They read the logs to understand the spy’s gameplay situation when establishing initial labels.
-To avoid bias towards a specific LLM, we hid the information of the models in the process.
+To avoid bias towards a specific LLM, we hid the information about the models in the process.
 Second, we refined these initial labels through an iterative process.
@@
 We repeated this process five times between December 29, 2023 and June 1, 2024 until agreement on the identified labels was saturated.
-The agreement was measured using Fleiss’ Kappa[24], and the final agreement between the annotators ranged from 0.486 to 0.646, which indicates moderate agreement.
+The agreement was measured using Fleiss’ Kappa[27], and the final agreement between the annotators ranged from 0.486 to 0.646, which indicates moderate agreement[28].
 Finally, using these settled labels, we labeled 506 reasoning steps from 168 games.
 Consequently, we identified three major categories and five subcategories.
-TABLE IV: Frequency of each category found in qualitative analysis
-$N$* Character ambiguity Memory distortion Dissociation
+TABLE VIII: Frequency of each category found in qualitative analysis
+NN* Character ambiguity Memory distortion Dissociation
 Against weak citizens
@@
 For each category, we explain its definition, report its frequency, and provide some examples.
-Table IV summarizes the frequencies of all categories.
+Table VIII summarizes the frequencies of all categories.
 V-B1 Exposure
@@
 [Player4, the spy, tries to guess the hidden location.]
-Llama2 (Spy): Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...
-We suspect that this category occurs because LLMs are not good at differentiating between numerical symbols.
+Llama2 (Spy): ‘Based on the conversation, I would say that the location is most likely a workplace or office. Player4 described their environment as well-lit, well-ventilated, ...’
+We suspect that this category occurs because LLMs struggle to differentiate between numerical symbols.
 As previous studies have reported that LLMs suffer from biases induced by proper names[29], we named each player as ‘Player#’ (e.g., Player1) to avoid such biases.
@@
 [Previously, Player2 and Player3 did not have any conversation. The spy tries to guess the hidden location.]
-GPT-3.5 (Spy): Based on the conversation between Player2 and Player3, it seems that the location is likely ...
+GPT-3.5 (Spy): ‘Based on the conversation between Player2 and Player3, it seems that the location is likely ...’
 We suspect this category is identical to the common “hallucination” phenomenon of LLMs, where a language model generates non-sensical or unfaithful text for the provided source input[31].
@@
 These two subcategories can occur independently.
-We present details of the subcategories in the below paragraphs.
+We present details of the subcategories in the following paragraphs.
 In summary, Fleiss’ Kappa for this category was 0.563.
@@
 [The spy tries to answer a citizen’s question. Before that, the spy generates its reasoning first.]
-Gemini (Spy): Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...
+Gemini (Spy): ‘Emphasize the importance of ambiance and personalized service, while avoiding giving away any specific details about the location...’
 Goal misunderstanding
@@
 [The spy tries to decide the next action. After selecting ‘question’ as the next action, the spy generates its reasoning.]
-Llama2 (Spy): ... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...
-We suspect that this category occurs is due to the memory limitation of LLMs.
+Llama2 (Spy): ‘... By asking this question, I hope to gain more information about the location and potentially identify the spy. If the spy is unaware of the restaurant’s cuisine, ...’
+We suspect that this category occurs due to the memory limitation of LLMs.
 To understand a player’s identity correctly, an LLM should retain information about the player’s identity during the decoding process.
@@
 However, such a distinction is difficult for the LLM because the other parts of the input prompt are much larger than the identity part.
-For example, the explanation of the game rules takes more than 130 words, whereas that of identity takes only 10 words.
-Thus, LLM players can confuse their identities with other characters written in the game rules, triggering character ambiguity.
-V-C Reconfirmation of Quantitative Analysis
-Combining the results of the two analyses confirms the findings of the quantitative analysis.
-Specifically, we discovered that these quantitative labels, except exposure, can be connected to two abilities in the quantitative analysis: subtext inference and deceptive control.
+For example, the explanation of the game rules holds more than 130 words, whereas that of identity takes only 10 words.
+Thus, LLM players can confuse their identities with other characters described in the game rules, triggering character ambiguity.
+V-C Discussion
+In this section, we first discuss how the reasoning error categories identified via thematic analysis help interpret the performance differences in subtext inference and deceptive control observed in our quantitative metrics.
+Then, we extend this discussion by exploring the potential generalizability of these error categories to broader LLM applications beyond the game-based setting.
+TABLE IX: Linking reasoning failures to quantitative performance
+Core Aspect Reasoning Error Behavioral or Reasoning Effect Quantitative Effect
+Subtext Inference Memory Distortion & Dissociation Incorrect deduction of hidden location, leading to failure of Guessing action lower GSR
+lower ICR
+lower IDR
+Deceptive Control Character Ambiguity Increasing suspicion from other players, leading to failure of misleading opponents higher CR
+higher VR
+lower VE
+As illustrated in Table IX, we map the identified reasoning errors to their quantitative impacts on the reasoning processes or behaviors of LLMs during the game.
 First, memory distortion and dissociation may hinder a spy from inferring subtext during the game.
-When we compared the frequency of memory distortion and dissociation with the results of the four subtext inference metrics, the trend of frequency was similar to that of the metrics.
-For example, against strong citizens, GPT-3.5 showed the highest errors in both memory distortion and dissociation and achieved the lowest performance in subtext inference.
-Similarly, GPT-4 showed fewer errors while achieving a higher performance than the other models.
-We suspect that this similarity between trends is due to the quality of information that the spy considers during their reasoning process.
-The spy player should select and combine pieces of information from the conversation to identify the hidden location.
-When memory distortion occurs, the spy can regard misinterpreted pieces as appropriate.
-Similarly, when dissociation occurs, the spy regards apocryphal information that they previously said as information provided by citizens.
-Therefore, such misunderstood pieces can be obstacles to reaching the correct answer.
+When we compared the frequency of memory distortion and dissociation with the results of the four subtext inference metrics, the frequency trend was similar to that of the metrics.
+For example, against strong citizens, GPT-3.5 exhibited the highest number of memory distortion and dissociation errors, and also showed the lowest performance across the subtext inference metrics.
+Similarly, GPT-4 exhibited fewer memory distortion and dissociation errors, along with higher subtext inference performance compared to the other models.
+We interpret this alignment as stemming from the quality of information the model considers during reasoning.
+To infer the hidden location, the spy should capture or synthesize clues from the conversation.
+However, when memory distortion occurs, the spy may misremember or misinterpret prior utterances as valid sources.
+Likewise, in the case of dissociation, the model may mistakenly treat its own prior statements as valid information provided by others.
+These misunderstandings can lead to incorrect conclusions, hindering the spy’s successful inference of the subtext.
 Second, character ambiguity increases the likelihood of a spy’s failure of deceptive control.
-When we compared the frequency of character ambiguity and the results of the three metrics regarding deceptive control, the trend of frequency was similar to that of the metrics.
+When we compared the frequency of character ambiguity and the results of the three metrics regarding deceptive control, the frequency trend was similar to that of the metrics.
 For example, against strong citizens, GPT-3.5 showed the lowest errors in character ambiguity and achieved the highest performance in such metrics.
@@
 Such confusion leads the spy to make an unusual utterance, which draws suspicion.
+Further, these findings not only highlight key failure modes in obscured communication but also point to a broader applicability of these reasoning errors across diverse LLM use cases.
+While our experimental setup is based on a social deduction game, many components of our qualitative analysis are not inherently limited to game-based settings.
+For example, memory distortion can affect retrieval-augmented generation (RAG) or multi-turn QA, in which LLMs fail to retain retrieved or prior context appropriately.
+Likewise, character ambiguity is also relevant in role-specific instruction following, such as persona-based assistants or simulation environments, where confusion about LLMs’ intended role can result in inconsistent or constraint-violating responses.
+These observations suggest that the taxonomy of reasoning failures developed in this study may offer a useful framework for diagnosing and improving LLM behavior beyond the context of obscured communication.
 VI Conclusion
-To assess whether LLMs can effectively support obscured communication, we addressed two key limitations in prior evaluation methodologies.
-The first is the reliance on coarse-grained quantitative metrics, which fail to capture fine-grained communicative behaviors.
-The second is the lack of structured qualitative analyses that can contextualize evaluation results or uncover the reasoning processes underlying model behavior.
-To investigate these limitations in a controlled setting, we conducted experiments using SpyGame, a variant of Spyfall, and collected gameplay logs from four LLMs.
-We analyzed these logs through both quantitative and qualitative methods.
-First, we proposed eight fine-grained metrics designed to evaluate two core abilities: subtext inference and deceptive control.
-These metrics demonstrated more effective in identifying subtle behavioral differences between models.
-Second, we performed a thematic analysis to identify reasoning failures not captured by quantitative metrics.
-This analysis revealed four major categories of failure that align with the patterns observed in the quantitative evaluation, offering complementary insights.
-Our findings highlight the importance of microscopic, multi-perspective evaluation for understanding how LLMs perform in obscured communicative settings.
-We hope this study contributes to a deeper understanding on LLMs in complex social interactions.
+In this study, we formulated two research questions to guide our investigation into how LLMs support obscured communication.
+The first concern is whether prior coarse-grained outcome metrics are sufficient for evaluating the fine-grained skills required in such contexts.
+The second investigates the underlying reasoning failures that can be systematically identified, and how these failures hinder LLMs’ ability to perform effectively in adversarial communicative scenarios.
+To examine the first question, we proposed six fine-grained quantitative metrics targeting two core dimensions of obscured communication: subtext inference and deceptive control.
+Through empirical evaluation using gameplay data collected from four LLMs in the SpyGame environment, we demonstrated that these metrics facilitate the diagnosis of intermediate behavioral patterns and expose the limitations of conventional evaluation methods that rely solely on outcome-level indicators such as win rate.
+To address the second question, we conducted a thematic qualitative analysis of gameplay logs capturing the LLMs’ reasoning processes and actions.
+This analysis revealed recurring reasoning error categories, such as memory distortion, dissociation, and character ambiguity, which not only reflect systematic failure modes but also align with patterns observed in the quantitative results.
+This alignment provides explanatory insight into the underlying causes of LLM performance breakdowns in obscured communication tasks.
+Despite our contribution, we aim to broaden its scope in future work by improving two points.
+First, we plan to extend and validate our analysis beyond the current game-based environment, exploring its generalizability to broader domains.
+In parallel, we will expand the evaluated roles beyond the single-agent spy to include diverse multi-agent settings, enabling the assessment of collaborative deception and coordination skills that are difficult to capture in the current setup.
+Overall, this study contributes a structured evaluation framework that combines fine-grained metrics with qualitative error analysis, offering a deeper understanding of LLM behavior in complex communicative settings.
+We hope this study contributes to a deeper understanding of LLMs in complex social interactions.